### v1

# Compatible Gradient Approximations for Actor-Critic Algorithms

Abstract
Deterministic policy gradient algorithms are foundational for actor-critic methods in controlling continuous systems, yet they often encounter inaccuracies due to their dependence on the derivative of the critic’s value estimates with respect to input actions. This reliance requires precise action-value gradient computations, a task that proves challenging under function approximation. We introduce an actor-critic algorithm that bypasses the need for such precision by employing a zeroth-order approximation of the action-value gradient through two-point stochastic gradient estimation within the action space. This approach provably and effectively addresses compatibility issues inherent in deterministic policy gradient schemes. Empirical results further demonstrate that our algorithm not only matches but frequently exceeds the performance of current state-of-the-art methods.
1 Introduction
Reinforcement learning (RL) has established itself as a prominent and effective method for addressing dynamic decision problems [32]. A fundamental strategy in tackling RL problems is via policy gradient (PG) techniques [33], grounded in the hypothesis that actions are selected according to a parameterized distribution. PG methods iteratively update policies via stochastic gradient schemes [33], enabling solutions in complex problems featuring continuous state-action spaces under uncertainty. PG methods are often framed within an actor-critic framework [11], where an actor determines the control policy evaluated by a critic. Thus, the efficacy of the actor’s improvements critically depends on the quality of the feedback provided by the critic. This interdependence raises nontrivial questions on the efficiency of training strategies for both the actor and the critic [5].
Conventionally, the critic is optimized through temporal difference (TD) learning [31] to accurately predict the expected returns led by the actor’s decisions. If the control policy is deterministic, the value estimated by the critic is solely used to compute the action-value gradient during policy optimization [5]. This involves computing the gradient of the critic’s action-value estimates with respect to the actions selected by the policy. More precisely, the deterministic policy gradient (DPG) [29] is determined by chaining the actor’s Jacobian to the action-gradient of the critic. However, TD learning optimizes a critic that learns the action-value, rather than its gradient. Generally, when function approximation is used for the critic in DPG, the resulting PG may not align with the true gradient. In some cases, it may not even constitute an ascent direction [29]. Nevertheless, this failure mode can be circumvented provided that $Q$-function estimates are compatible, meaning a set of conditions necessary for (approximately) following the true gradient; see, e.g., Theorem 3 of Silver et al. [29].
Integration of high-dimensional scenarios requires complex function approximations, such as deep neural networks [13, 7, 8]. However, the use of neural networks violates the classical compatibility results, which (conveniently) calls for critics with linear gradient representations [29, Theorem 3]. Hence, contemporary sample-based algorithms contradict the very foundations of DPG (from the angle of compatibility), thereby paving the way for potential failure scenarios (see Appendix A.3 for an illustration). Indeed, action differentiation of the $Q$-function (estimates), i.e., action-value gradients, leaves modern successors to DPG technically unsound [5].
In this paper, we introduce (off-policy) Compatible Policy Gradient (oCPG), a continuous control actor-critic algorithm that approximates the action-value gradient using only value estimates from the critic. The proposed method embodies a zeroth-order nature by removing the direct computation of the action-value gradient, while provably addressing the compatibility requirement in DPG. This is achieved through evaluating critics at low-dimensional random perturbations within the action space, resulting in (batch) two-point policy gradient approximations whose distance to the true deterministic policy gradient is controllable at will.
The performance of oCPG is assessed on challenging OpenAI Gym continuous control tasks [4], in perfect (less uncertain) and imperfect (higher uncertainty) environmental conditions. Our numerical results demonstrate that oCPG exhibits robust performance and either matches or surpasses (often substantially) the state-of-the-art in terms of stability and mean rewards. For reproducibility purposes, we share our code in the GitHub repository [^1] https://github.com/baturaysaglam/compat-grad-approx.
2 Related Work
Policy gradient techniques are widely used in RL. Various algorithms for estimating policy gradients have been developed; Baxter and Bartlett [3], Metelli et al. [16], Williams [35] focus solely on the policy, while Mnih et al. [17], Schulman et al. [27, 28], Konda and Tsitsiklis [11], Prokhorov and Wunsch [20] also incorporate a value function, known as actor-critic methods. Particularly, algorithms based on the deterministic policy gradient [13, 29] rely on the action-gradient from the critic. In function approximation, the accuracy of the critic is crucial [2]; for example, applying compatibility conditions to the critic helps achieve an unbiased estimate of the policy gradient [29].
The incompatibility of the critic in DPG could be resolved by employing zeroth-order optimization (ZOO), which replaces the action-value gradient computation with an approach that does not require any (first-order) gradient information. These techniques treat the RL problem as a black-box optimization by ignoring the underlying MDP structures and directly searching for the optimal policy without gradient computations. Recent ZOO methods have proven to be competitive on common RL benchmarks [15, 24, 12], benefiting especially from not being constrained to differentiable policies. Despite these advantages, ZOO techniques suffer from high sample complexity and significant variance in gradient updates, due to their disregard for the MDP structure [12]. These issues are further amplified with the increase in perturbation noise scale, potentially escalating dramatically as parameter counts can reach into the millions [6, 12]. Moreover, while parameter space perturbations can induce diverse behaviors viewed as exploration, they are less beneficial in DPG methods, which are typically framed in an off-policy context [29, 13, 7]. Exploration and policy learning are distinct processes in off-policy methods, making parameter-based zeroth-order techniques less effective for enhancing exploration, primarily confined to on-policy and stochastic policies [30]. Contrasting with parameter space-based ZOO techniques, our approach operates in the action space and involves low-dimensional perturbations. Perturbing the action also incorporates Gaussian exploration [7], which has been empirically shown to outperform parameter space exploration in off-policy settings [22, 21]. Lastly, our objective is not merely to devise a pure zeroth-order PG method but to synthesize the strengths of both ZOO and first-order approaches, employing a differentiable policy whose Jacobian is chained to the zeroth-order approximation of the action-value gradient.
In addition to the zeroth-order perspective, GProp [1] and MAGE [5] are the closest known methods to our study that directly focus on optimizing and ensuring accuracy in the action-value gradient computation. GProp modifies traditional TD learning [31] by using an additional neural network that predicts the action-value gradient. MAGE, drawing on recent theoretical [25] and practical [26] insights, trains an additional network to simulate environmental dynamics, employing it as a proxy to sample states for double gradient descent on the $Q$-function. In stark contrast, our method (CPG) simplifies the DPG approach by directly estimating the action-value through a zeroth-order method, employing two-point evaluations of the existing critic network. This eliminates the need to train additional networks, modify TD learning frameworks, or depend on model-based settings. CPG is adaptable to a broader range of problems, bypassing the requirements for double differentiability of the $Q$-function and detailed modeling of transition dynamics. Additionally, our approach seamlessly integrates with any TD learning method and policy learning framework, and requires only a few lines of code to adapt to modern DPG-based algorithms.
3 Technical Preliminaries
We consider a standardized RL problem in which an agent moves through an environment characterized by a continuous compact state space $\mathcal{S}\subset\mathbb{R}^{q}$ and takes actions in a finite dimensional action space $\mathcal{A}=\mathbb{R}^{p}$. Based on its action selection, the agent receives a reward from a (bounded, for simplicity) reward function $R$, where $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and observes the next state $s^{\prime}\in\mathcal{S}$. This generic problem is often abstracted by a Markov decision process (MDP) as a tuple $(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $P$ denotes the environment dynamics, that is, the probabilities $p(s^{\prime}|s,a)$ of moving to state $s^{\prime}$ from state $s$ and if action $a$ performed, satisfying the Markov property: $p(s_{t+1}|s_{i},a_{i};\forall i\leq t)=p(s_{t+1}|s_{t},a_{t})$. The value $\gamma\in(0,1)$ is the discount factor that prioritizes future rewards.
The action behavior of the agent is described by a policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$ parameterized by $\theta\in\Theta$, which can be either a deterministic function or represent a probability distribution. Note that we may occasionally omit the parameter subscript. The agent’s objective is to maximize the expected discounted sum of rewards, as quantified by the value function: $V^{\pi_{\theta}}(s)\coloneqq\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t}\sim\pi_{\theta}(\cdot\mid s_{t}))|s_{0}=s]$. By conditioning on an action, we may also define the action-value function $Q^{\pi_{\theta}}(s,a)\coloneqq\mathbb{E}\left[R(s_{0},a_{0})+\sum_{t=1}^{\infty}\gamma^{t}R(s_{t},a_{t}\sim\pi_{\theta}(\cdot\mid s_{t}))|s_{0}=s,a_{0}=a\right]$. Lastly, let $\rho^{0}$ be the initial state distribution, and let $\rho^{\pi}$ be the $\gamma$-discounted state probability distribution induced by policy $\pi$, defined as $\rho^{\pi}(s^{\prime})=\int_{\mathcal{S}}(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}p_{0}(s)p(s\rightarrow s^{\prime},t,\theta)\mathop{}\!\mathrm{d}s$, where $p(s\rightarrow s^{\prime},t,\theta)$ denotes the probability density (we assume such a density exists for simplicity) at state $s^{\prime}$ after starting from state $s$ and transitioning for $t$ time steps under a policy parameterized by $\theta$.
At the initial state $s$, by selecting the first action according to the policy and advancing the system, the $Q$-function becomes equivalent to the value function. Consequently, the objective is expressed as
$J(\theta)\coloneqq\underset{\theta\in\Theta}{\max}\text{\phantom{ }}\mathbb{E}_{s\sim\rho^{0},a\sim\pi_{\theta}(\cdot|s)}\bigr{[}Q^{\pi_{\theta}}(s,a)\bigr{]}.$
When the search is confined to deterministic policy parameterizations (and we do so hereafter), the problem above reduces to a simpler form, i.e.,
$J(\theta)=\underset{\pi\in\Pi}{\text{max}}\text{\phantom{ }}\mathbb{E}_{s\sim\rho^{0}}\bigr{[}Q^{\pi_{\theta}}(s,\pi_{\theta}(s))\bigr{]}.$ (1)
Under mild problem regularity conditions, it has been demonstrated by Silver et al. [29] that the gradient of a deterministic policy may be expressed by
$\nabla J(\theta)=\frac{1}{1-\gamma}\int_{\mathcal{S}}\rho^{\pi_{\theta}}(s)\nabla_{\theta}\pi_{\theta}(s)\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{|}_{a=\pi_{\theta}(s)}\mathop{}\!\mathrm{d}s$
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\pi_{\theta}}}\bigr{[}\nabla_{\theta}\pi_{\theta}(s)\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{|}_{a=\pi_{\theta}(s)}\bigr{]},$ (2)
where we use the normalizing term $1/(1-\gamma)$ since $\rho^{\pi_{\theta}}$ is proper. The latter equation is known as the deterministic policy gradient theorem [29], which has been instrumental in the development state-of-the-art policy gradient methods essentially relying on stochastic approximation, such as Deep Deterministic Policy Gradient (DDPG) [13] and Twin Delayed DDPG (TD3) [7].
4 Off-Policy Compatible Policy Gradient
4.1 Problem Statement
Using a function-approximated $Q$-function in the DPG algorithm, as expressed in (2), can disrupt the accuracy of the approximated PG unless specific conditions are met [29, Theorem 3]. For completeness, we provide a detailed summary of these compatibility requirements in Appendix A.2. A critical point of Silver et al. [29] is that the gradient of the $Q$-function is conveniently assumed to be linear, i.e., $\nabla_{a}Q_{\omega}(s,a)|_{a=\pi_{\theta}(s)}=\nabla_{\theta}\pi_{\theta}(s)^{\top}\omega$ for parameters $\omega$. However, it has been demonstrated that scaling to large and complex state-action spaces inevitably requires the use of more expressive representations, such as deep neural networks [13]. This necessity compromises compatibility and results in policy gradients that may not accurately follow the true gradient [29].
This failure mode has recently been linked to the need to minimize the norm of the action-value gradient of the policy evaluation error, rather than its value, in order to reduce the approximation error of the computed PG [5, Proposition 3.1]. Aligning with this insight, double backpropagation through the TD error may address this issue in scenarios where model-based approaches are feasible [5]. However, model-free RL is often preferred due to its effectiveness in environments with complex or unknown dynamics, offering enhanced flexibility and robustness for direct policy learning via interactions with the environment [32].
While several studies have pursued the goal of overcoming the incompatibility of the critic [5, 1], our goal herein is to consistently approximate the deterministic policy gradient $\nabla J$ as
^$\hat{\nabla}J(\theta)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\pi}}\big{[}\nabla_{\theta}\pi_{\theta}(s)\hat{\nabla}_{a}Q^{\pi_{\theta}}(s,a))\big{|}_{a=\pi_{\theta}(s)}\big{]},$ (3)
where ^$\hat{\nabla}_{a}Q^{\pi_{\theta}}(s,a)$ represents an appropriately designed approximation of ${\nabla}_{a}Q^{\pi_{\theta}}(s,a)$, resulting in a ^$\hat{\nabla}J$ that is provably compatible, general and implementable in a model-free manner. To the best of our knowledge, no study has yet proposed a gradient surrogate (i.e., ^$\hat{\nabla}J$) that possesses all aforementioned features simultaneously, in a model-free deep RL setting. As we empirically demonstrate later (Section 5), such an approach in fact results in substantial operational benefits over the state-of-the-art.
4.2 Provably Compatible Policy Gradient Approximations
To begin, we temporarily assume that $Q^{\pi_{\theta}}$ is available and, for a smoothing parameter $\mu>0$, we consider the $\mu$-smoothed $Q$-function defined as
$Q^{\pi_{\theta}}_{\mu}(s,a)\coloneqq\mathbb{E}_{\boldsymbol{u}}\big{[}Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})\big{]},\,\,\boldsymbol{u}\sim\mathcal{N}(0,I_{p}).$ (4)
Smoothing may be equivalently thought of as enforcing exploration, and the random perturbation $\mu\boldsymbol{u}$ may be thought of as the stochastic part of a standard Gaussian policy $\pi_{\theta}^{\mu}(\cdot\mid s)\coloneqq\pi_{\theta}(s)+\mu\boldsymbol{u}$, with $\boldsymbol{u}\sim\mathcal{N}(0,I_{p})$. Here, it will be useful to think of such a Gaussian policy as a deterministic policy perturbed by leveraging low-dimensional noise in action space.
Implementing (Gaussian) smoothing on deterministic actions aligns with approaches used in seminal works such as by Lillicrap et al. [13], Fujimoto et al. [7]. As mentioned above, this technique introduces stochasticity in action selection, which prevents premature convergence to suboptimal policies by promoting exploration of a broader range of state-action pairs [32, 7]. In any case, though, the ultimate goal is to discover a near-optimal deterministic control policy that near-solves (1). In our context, we achieve this by learning a deterministic policy through trajectories generated by a smoothed/perturbed policy, with all subsequent evaluations based on deterministic actions.
Our development will be relying on the following basic result by Nesterov and Spokoiny [19], which is central in the development and analysis of zeroth-order optimization methods, and establishes that the gradient of the smoothed functions such as that in (4) may be evaluated in a model-free fashion, through batches of two-point evaluations of the function itself.
Proposition 1 [19].
Let $f:\mathbb{R}^{p}\rightarrow\mathbb{R}$ be a bounded function. For every $\mu>0$, the smoothed function $f_{\mu}(x)\coloneqq\mathbb{E}_{\boldsymbol{u}}\big{[}f(x+\mu\boldsymbol{u})\big{]}$, $\boldsymbol{u}\sim\mathcal{N}(0,I_{p})$ is well-defined, differentiable and its gradient admits the representation
$\nabla f_{\mu}(x)=\mathbb{E}_{\boldsymbol{u}\sim\mathcal{N}(0,I_{n})}\Bigg{[}\frac{f(x+\mu\boldsymbol{u})-f(x)}{\mu}\boldsymbol{u}\Bigg{]},\,\,\forall x\in\mathbb{R}^{p}.$
Further, if $f$ is $G$-smooth (i.e., with $G$-Lipschitz gradients), it holds that
${\sup_{x}}\|\nabla f_{\mu}(x)-\nabla f(x)\|\leq\mu G\sqrt{p}.$
Proposition 1 easily motivates the $Q$-function zeroth-order gradient approximation
${\nabla}_{a}Q_{\mu}^{\pi_{\theta}}(s,a)=\mathbb{E}_{\boldsymbol{u}}\Bigg{[}\frac{Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a)}{\mu}\boldsymbol{u}\Bigg{]},$
resulting in a deterministic policy gradient approximation reading (cf. 3)
^$\hat{\nabla}^{\mu}J(\theta)\coloneqq\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\pi_{\theta}},\boldsymbol{u}\sim\mathcal{N}(0,I_{p})}\Bigg{[}\nabla_{\theta}\pi_{\theta}(s)\times\frac{Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a)}{\mu}\boldsymbol{u}\Bigg{|}_{a=\pi_{\theta}(s)}\Bigg{]}.$
While ^$\hat{\nabla}^{\mu}J$ is a genuine and consistent model-free approximation of the deterministic policy gradient ${\nabla}J$ (through Proposition 1), it requires access to (evaluations of) the $Q$-function itself, the latter being both unknown (in the majority of cases) and also hard, or at least non-trivial to sample.
Approximating the $Q$-function
Following the actor-critic paradigm [11], we approximate the $Q$-function by a parameterized learning representation $Q_{\psi}$ (the critic), for instance, a deep neural network (i.e., $Q$-network), similarly to the control policy $\pi_{\theta}$ (or actor). We denote $Q$-network parameters abstractly with the subscript $\psi$ (different than the smoothing parameter $\mu$). Then, we propose the parameterized deterministic policy gradient approximation
^$\hat{\nabla}^{\mu,\psi}J(\theta)\coloneqq\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\pi_{\theta}},\boldsymbol{u}}\Bigg{[}\nabla_{\theta}\pi_{\theta}(s)\times\frac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\Bigg{|}_{a=\pi_{\theta}(s)}\Bigg{]}.$ (5)
Given (5), a natural question is how the representation $Q_{\psi}$ should be trained, in order for ^$\hat{\nabla}^{\mu,\psi}J$ to achieve a small approximation error as compared with the true policy gradient $\nabla J$. In other words, instead of approximating in value, we would like to choose $\psi$ such that the representation $Q_{\psi}$ is compatible to $Q^{\pi_{\theta}}$ (in the standard sense of Silver et al.). To this end, let us first define the perturbed representation error
$\varepsilon_{\mu,\psi}^{\pi_{\theta}}\coloneqq\sqrt{\mathbb{E}_{s,\boldsymbol{u}}\big{[}\big{|}Q_{\psi}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})\big{|}^{2}\big{]}},$
where $\theta,\mu$ and $\psi$ are given. We have the following result (see Appendix A.4 for the proof).
Theorem 1 Compatible Policy Gradient.
Let $B>0$ be a constant such that $\sup_{(s,\theta)\in\mathcal{S}\times\Theta}\|\nabla_{\theta}\pi_{\theta}(s)\|\leq B$. Then the gradient approximation ^$\hat{\nabla}^{\mu,\psi}J$ satisfies, for every $\theta\in\Theta,\mu>0$ and $\psi$,
^$\|\hat{\nabla}^{\mu,\psi}J(\theta)-\nabla J(\theta)\|\leq\dfrac{B}{1-\gamma}\bigg{[}\dfrac{\varepsilon_{\mu,\psi}^{\pi_{\theta}}}{\mu}\sqrt{p}+\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]}\bigg{]}.$
In particular, if $Q^{\pi_{\theta}}(s,\cdot)$ is itself $G$-smooth (uniformly), it follows that, for every $\theta\in\Theta$,
^$\|\hat{\nabla}^{\mu,\psi}J(\theta)-\nabla J(\theta)\|\leq\dfrac{B\sqrt{p}}{1-\gamma}\bigg{[}\dfrac{\varepsilon_{\mu,\psi}^{\pi_{\theta}}}{\mu}+G\mu\bigg{]}.$
The usefulness of the result just stated is twofold: One the one hand, Theorem 1 cleanly quantifies the performance of ^$\hat{\nabla}^{\mu,\psi}J$ in terms of smoothing ($\mu$) and function approximation ($\psi$) errors. On the other hand, Theorem 1 reveals how $\psi$ can be chosen, so that the gradient approximation error achieved by adopting ^$\hat{\nabla}^{\mu,\psi}J$ is controlled at will. To see this, observe that if we take for clarity the case of a smooth $Q$-function, Theorem 1 implies that
$\inf_{\psi}\|\hat{\nabla}^{\mu,\psi}J(\theta)-\nabla J(\theta)\|\leq\dfrac{B\sqrt{p}}{1-\gamma}\bigg{[}\dfrac{{\textstyle\inf_{\psi}\varepsilon_{\mu,\psi}^{\pi_{\theta}}}}{\mu}+G\mu\bigg{]}.$
In other words, for a fixed level of smoothing $\mu>0$ (and at each actor instance indexed by $\theta$), one can optimally choose (i.e., train) the representation $Q_{\psi}$ such that the perturbed representation error $\varepsilon_{\mu,\psi}^{\pi_{\theta}}$ is made as small as possible, theoretically achieving an arbitrarily small gradient approximation error (by properly choosing an expressive enough class for the representation $Q_{\psi}$). Motivated from the discussion above, we refer to ^$\hat{\nabla}^{\mu,\psi}J$ as a Compatible Policy Gradient (CPG).
Discussion
In passing, we would like to emphasize that although Gaussian exploration noise appears in $\varepsilon_{\mu,\psi}^{\pi_{\theta}}$, the ultimate goal is deterministic policy optimization, as seen in off-policy methods such as DDPG [13, Algorithm 1, lines 8-13] and TD3 [7, Algorithm 1, lines 6-11]. However, both DDPG and TD3 use the gradient of the fitted $Q$-function. Theorem 1 suggest that (5) should instead be used as the (provably compatible) gradient approximation of choice. The reason is that sample-based first-order PG methods use a finite number of samples (i.e., collected transition tuples) to fit their action-value functions. However, accurately approximating function values does not necessarily imply accurate gradient approximation. In fact, for any finite number of samples, one can fit a function with zero approximation error (say, in the mean-square sense), but with arbitrarily distinct gradients at each point, as compared with the original function. In Appendix A.3, we discuss this through an intuitive illustration. Therefore, it is important to note that the main goal of our study is to ensure that the PG is directly connected to the values of the fitted $Q$-function (using finite samples), instead of the gradients of the fitted $Q$-function. Lastly, recent studies (e.g., Saglam and Kozat) consider learnable exploration noise, where $\boldsymbol{u}$ is typically the output of a neural network. We discuss the applicability and potential outcomes of these methods in Appendix A.5.
4.3 Off-Policy Deep Reinforcement Learning
To devise an efficient learning framework centered on CPG, we incorporate standard components from existing literature.
Clipped Double $Q$-learning [7]
Based on our discussion above, to exploit ^$\hat{\nabla}^{\mu,\psi}J$ as a gradient surrogate (and thus harness the benefits of Theorem 1), we should learn a $Q_{\psi^{*}}$ such that $\psi^{*}\in\arg\min_{\psi}\varepsilon_{\mu,\psi}^{\pi_{\theta}}$ (iteration-wise for fixed $\theta$ and $\mu$). However, learning such a $Q_{\psi^{*}}$ assumes the feasibility of obtaining unbiased estimates of the $Q$-function at any state-action (perturbation) pair, as well as sampling from the discounted state distribution $\rho^{\pi}$. Specifically for obtaining unbiased $Q$-value estimates in the context of deep RL, one can employ the practical clipped double $Q$-learning algorithm [7] to learn the $Q$-network (in conjunction with an experience replay buffer – see next paragraph), which has been shown to eliminate the estimation bias effectively.
Off-Policy Learning
Our selection of an off-policy learning approach is driven by its potential for high sample efficiency, notably through the use of an experience replay buffer [14]. The agent stores samples as transition tuples $(s,a,R(s,a),s^{\prime})$ in the buffer and periodically samples a minibatch of these transitions for policy and $Q$-network updates. Building upon this, CPG, although theoretically on-policy, is implemented in an off-policy framework. This approach, however, inherently involves a distributional shift, as the samples used may significantly diverge from the current agent’s policy [32]. Despite this, our method draws on the strategy employed by modern off-policy algorithms [29, 13, 7, 8], which simplifies complexities by approximating off-policy samples as on-policy.
Combining the components introduced, we present our proposed CPG method in an off-policy deep RL setting, implementing stochastic gradient ascent via the implementable gradient approximation
^$\boxed{\hat{\nabla}^{\mu,\psi}J(\theta)\approx\hat{\nabla}_{\beta}^{\mu,\psi}J(\theta)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\beta},\boldsymbol{u}}\Bigg{[}\nabla_{\theta}\pi_{\theta}(s)\frac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\Bigg{|}_{a=\pi_{\theta}(s)}\Bigg{]},}$
where $\beta$ denotes the set of policies that collected the off-policy samples (referring to the replay buffer), which can also be on-policy. Acknowledging the limitations of this approximation, we focus on the immediate performance of our algorithm. We note that a detailed analysis of the “off-policyness” (see, e.g., Munos et al., Saglam et al.), as well as the complexity and convergence analysis of this actor-critic setup under CPG, are topics of our current scope, and thus deferred for future research.
Pseudocode for off-policy CPG (oCPG) is provided in Algorithm 1 in Appendix A.1. The inclusion of dual $Q$-networks, target parameters, and delayed policy updates draws from the structure of Clipped Double $Q$-learning. We also consistently optimize the policy with respect to the first $Q$-network to ensure stability.
5 Experiments
We evaluate the performance of oCPG by benchmarking it against state-of-the-art PG-based actor-critic algorithms: TD3 [7] and Soft Actor-Critic (SAC) [8]. Notably, as TD3 is a variant of DPG, comparing it with CPG serves as an effective test of our proposed PG method against the traditional DPG. Similarly, since SAC is an extension of the standard stochastic PG algorithm, our comparisons with SAC offer insights into CPG relative to stochastic PG.
5.1 Experimental Setup
Evaluation
The performance is assessed on a set of challenging continuous control tasks from the MuJoCo suite [34], interfaced by OpenAI Gym [4]. Each algorithm is run for 1 million time steps, except for Hopper, Humanoid, and Walker2d, which are extended to more steps. This adjustment is made because we observed a continuing upward trend in the rewards at 1 million time steps, indicating that more time was needed for the algorithms to reach a steady state. The agent is evaluated in a separate evaluation environment without exploration noise, i.e., the system is controlled by deterministic actions. Each evaluation occurs every 1000 time step for 10 episodes, and the average reward is recorded. Results are reported over 10 random seeds, including the Gym simulator, network initialization, and any other stochastic processes involved.
Implementation and hyperparameters
For SAC, we adhere to the hyperparameter setup as described by Haarnoja et al. [8]. The parameters for TD3 are imported from the authors’ GitHub repository [^2] https://github.com/sfujim/TD3, which provides the tuned setting. To mitigate dependency on initial policy parameters, we implement a purely exploratory policy during the first 25,000 time steps in all environments, a strategy shown to enhance sample complexity [7]. In the implementation of CPG, we omitted the $1-\gamma$ term in the denominator for simplicity. We used the same learning rate for a fair evaluation with TD3 to isolate the effect of our proposed PG method, as only the policy updates differ in this setting. Overall, we set the exploration degree to $\mu=0.1$ for all algorithms.
Uncertainty in the environments
To evaluate the robustness of oCPG, we simulate imperfect environmental conditions by introducing uncertainty. This includes considering scenarios with noisy, sparse, and delayed rewards. In the noisy reward scenario, each reward is subject to perturbation with random noise. For the sparse reward setup, the agent has a 0.5 probability of observing the actual reward; otherwise, it receives a reward of zero. In the delayed setting, the agent observes the reward after a fixed number of time steps.
A comprehensive description of the experimental setup and implementation is available in Appendix B.2
5.2 Results
The results are presented in Figure 1. Additionally, Table 1 provides the maximum rewards achieved, along with a Welch $t$-test to evaluate the statistical significance of the performance differences. Welch’s $t$-test is particularly suitable as it accounts for differences in variances between two independent groups, ensuring a robust comparison of their means. We also provide the results of different statistical tests, along with the learning curves under imperfect environmental conditions, in Appendix C.
TD3       SAC       oCPG
Figure 1: Learning curves for the benchmark MuJoCo environments, averaged over 10 random seeds. The shaded area represents the 95% confidence interval of the mean performance.
In most environments, oCPG achieves faster convergence to higher final rewards. Although the performance curves are intertwined, making it difficult to discern a clear leader visually, the statistical tests reported in Table 1 confirm that oCPG is the best-performing algorithm in three out of six environments. In the remaining cases, it is on par with the baselines.
Notice that the only difference between oCPG and TD3 is in how the PG is computed. The rest of the configurations, particularly the degree of exploration $\mu$ (which also influences the computed PG in our algorithm), remain identical. The statistically significant performance improvement is therefore attributable to the PG computation. While we do not directly claim it, this observation suggests that CPG may yield more accurate PG estimates than DPG.
Environment TD3 SAC oCPG
Ant 4225.080 $\pm$ 835.236 3945.508 $\pm$ 673.329 5284.923 $\pm$ 412.932
HalfCheetah 6744.592 $\pm$ 671.739 6074.230 $\pm$ 405.968 7956.502 $\pm$ 1178.024
Hopper 3620.488 $\pm$ 137.280 3585.355 $\pm$ 186.848 3687.673 $\pm$ 118.093
Humanoid 5289.623 $\pm$ 148.741 5567.597 $\pm$ 141.133 5348.386 $\pm$ 109.030
Swimmer 59.806 $\pm$ 9.477 58.396 $\pm$ 10.990 123.130 $\pm$ 59.253
Walker2d 4842.336 $\pm$ 723.696 4725.314 $\pm$ 492.717 4953.588 $\pm$ 298.321
Table 1: Average (best) performance on the benchmark MuJoCo environments over 10 random seeds, where $\pm$ denotes a 95% confidence interval. The highest performance, statistically superior to the others, is highlighted. Additionally, the highest performances shared by multiple algorithms are indicated, based on Welch’s $t$-test with a significance level of 0.05.
Moreover, while oCPG generally achieves faster convergence to higher rewards, we note slower convergence in the Humanoid environment compared to TD3 and SAC. The state-action dimension in Humanoid is aggregated to 393, whereas the $Q$-networks in our experiments have only 256 hidden units per layer. Neural networks can be sensitive to input noise, which is amplified when the input size exceeds the number of hidden units. This issue is likely intensified by our smoothing technique on line 13 in Algorithm 1, which introduces input noise to the neural network, potentially hindering performance in environments with large state-action spaces. However, this could be mitigated by using wider networks. Despite these challenges, oCPG still achieves higher maximum rewards, as confirmed in Table 1. Additionally, while the performance in the Swimmer task is statistically significant, the rewards display high variance. This variance stems from some seeds where oCPG converged to suboptimal levels, resulting in broad confidence intervals.
Discrepancies in reported reward levels as compared with those in the original TD3 and SAC studies may arise from variations in environmental stochasticity and the version of the MuJoCo simulator used. Nonetheless, consistent performance improvements over multiple random seeds provide a robust basis for assessing the efficacy of our method. This approach follows established deep RL experimentation standards [9], focusing on comparative analysis rather than absolute performance metrics.
Imperfect Environmental Conditions
We observe that while oCPG generally achieves the highest rewards or matches the baselines in worst-case scenarios, it exhibits a slower convergence rate in certain environments, such as Hopper and Walker2d. We attribute this to the nature of the reward-perturbing scenarios included in our tests. These scenarios significantly alter the rewards observed by the agent, disrupting the learning process of the $Q$-network where the reward is a direct input. Such changes impact the value estimates of the $Q$-network, which is directly used by CPG. Although the baselines also rely on the value estimate through its gradient, modifications to the rewards do not appear to significantly affect their final performance. This is further suggested by the observation that the precision of the gradient of the value estimates does not necessarily correspond to the precision of the value estimates themselves, as sketched in Appendix A.3. Despite these challenges, oCPG maintains robust performance across most environments.
Limitations
Our empirical results highlight slight limitations of oCPG. In large state-action spaces, the algorithm’s convergence speed may be hindered unless wider networks are employed. Similarly, convergence is adversely affected in environments with significantly altered rewards, such as imperfect settings, although oCPG can ultimately achieve higher reward levels over time. We bound the smoothing parameter $\mu$ to the exploration levels. Without this association, tuning $\mu$ for each environment is necessary. Still, our tests indicate that the fixed and unmodified value of $\mu=0.1$ can generalize across diverse environments, maintaining accurate policy gradient approximation and exploratory activity simultaneously.
Overall, the findings from our discussion indicate that CPG provides a robust and state-of-the-art alternative to traditional PG methods, especially in scenarios where the accuracy of action-value gradient computations is less reliable, effectively resolving the incompatibility issue.
6 Conclusions and Future Work
In this work, we have presented a policy gradient technique for actor-critic algorithms, Compatible Policy Gradient (CPG), which introduces a zeroth-order approximation method to estimate the action-value gradient in deterministic policy gradient (DPG) algorithms. Our approach circumvents the need for explicit action-value gradient computation by employing stochastic two-point gradient estimation by leveraging low-dimensional perturbations in action space. This effectively resolves compatibility issues often encountered in conventional DPG methods with function approximation. Empirical evaluations on various continuous control tasks from the OpenAI Gym suite demonstrate that CPG exhibits competitive performance as compared with the state-of-the-art methods.
The implications of our findings are twofold. First, they underscore the potential of zeroth-order (gradient-free) methods in overcoming the limitations of dependency on accurate derivative computations in policy optimization. Second, the efficacy of CPG in complex continuous control environments suggests that similar approaches could be beneficial in other domains of reinforcement learning where the dynamics are highly uncertain or the system models are imperfect.
Future Work
Developing an off-policy correction scheme for CPG could further improve its effectiveness. We believe that a scheme based on a ratio of the probabilities of the current agent and behavioral policies (collecting transitions) could provide an effective correction. Additionally, exploring the potential of the CPG approach in model-based RL scenarios could provide deeper insights into its versatility and robustness.
Appendix A Supporting Details for Section 4
A.1 Pseudocode of Off-Policy Compatible Policy Gradient
Algorithm 1 Off-Policy Compatible Policy Gradient (oCPG)
1:Initialize the policy network $\pi_{\theta}$, and critic networks $Q_{\psi_{1}}$, $Q_{\psi_{2}}$ with random parameters $\theta$, $\psi_{1}$, $\psi_{2}$
2:Initialize target networks $\theta^{\prime}\leftarrow\theta$, $\psi^{\prime}_{1}\leftarrow\psi_{1}$, $\psi^{\prime}_{2}\leftarrow\psi_{2}$
3:Initialize the experience replay buffer: $\mathcal{B}=\emptyset$
4:for $t=1$ to $T$ do
5:     Select action with Gaussian exploration noise $a=\pi_{\theta}(s)+\mu\boldsymbol{u}$, where $\boldsymbol{u}\sim\mathcal{N}(0,I_{p})$
6:     Execute action $a$, receive reward $r=R(s,a)$, and observe new state $s^{\prime}$
7:     Store the collected transition tuple in the replay buffer: $\mathcal{B}\leftarrow\mathcal{B}\cup(s,a,r,s^{\prime})$
8:     Sample a minibatch of $N$ transitions from the replay buffer: $(\mathbf{s},\mathbf{a},\mathbf{r},\mathbf{s}^{\prime})\sim\mathcal{B}$
9:     $\tilde{\mathbf{a}}\leftarrow\pi_{\theta^{\prime}}(\mathbf{s}^{\prime})+\epsilon$, where $\{\epsilon_{i}\sim\text{clip}\left(\mathcal{N}(0,\sigma^{2}I_{p}),-c,c\right)\}_{i=1}^{N}$triangleright Clipped Double $Q$-learning
10:     $\mathbf{y}\leftarrow\mathbf{r}+\gamma\text{ min}_{i=1,2}Q_{\psi^{\prime}_{i}}(\mathbf{s}^{\prime},\tilde{\mathbf{a}})$triangleright Clipped Double $Q$-learning
11:     Update critics: $\psi_{i}\leftarrow\text{argmin}_{\psi_{i}}\frac{1}{N}\sum_{i}\left(\mathbf{y}_{i}-Q_{\psi_{i}}(\mathbf{s}_{i},\mathbf{a}_{i})\right)^{2}$triangleright Clipped Double $Q$-learning
12:     if $t$ mod $d$ then $\triangleright$ Delayed policy updates
13:         Update the policy by CPG: ^${\hat{\nabla}_{\beta}^{\mu,\psi}J(\theta)=\dfrac{1}{N}\sum_{i}\nabla_{\theta}\pi_{\theta}(\mathbf{s}_{i})\dfrac{Q_{\psi_{1}}(\mathbf{s}_{i},\mathbf{a}_{i}+\mu\boldsymbol{u}_{i})-Q_{\psi_{1}}(\mathbf{s}_{i},\mathbf{a}_{i})}{\mu}\boldsymbol{u}_{i}\Bigg{|}_{\begin{subarray}{c}\mathbf{a}_{i}=\pi_{\theta}(\mathbf{s}_{i})\\ \boldsymbol{u}_{i}\sim\mathcal{N}(0,I_{p})\end{subarray}}}$
14:         Update target networks: $\theta^{\prime}\leftarrow\tau\cdot\theta+(1-\tau)\cdot\theta^{\prime}$ 15:$\psi_{i}^{\prime}\leftarrow\tau\cdot\psi_{i}+(1-\tau)\cdot\psi_{i}^{\prime}$
16:     end if
17:end for
A.2 Compatibility Requirements in Deterministic Policy Gradients
Here, for completeness we recall the classical result of Silver et al. [29] (specifically Theorem 3 in Silver et al. [29]), which ensures exact gradient compatibility, however at the expense of restrictive conditions on the form of the adopted $Q$-function approximation (in particular, its gradient must be of linear form).
Theorem 2 Compatible Function Approximation.
A function approximator $Q_{\omega}(s,a)$ is compatible with a deterministic policy $\pi_{\theta}(s)$, $\nabla J(\theta)=\mathbb{E}_{s\sim\rho^{\pi_{\theta}}}\bigr{[}\nabla_{\theta}\pi_{\theta}(s)\nabla_{a}Q_{\omega}(s,a)\big{|}_{a=\pi_{\theta}(s)}\bigr{]}$, if
  1. 1.
    $\nabla_{a}Q_{\omega}(s,a)\big{|}_{a=\pi_{\theta}(s)}=\nabla_{\theta}\pi_{\theta}(s)^{\top}\omega$ and
  2. 2.
    $\omega$ minimizes the gradient mean-squared error $\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\omega}(s,a)|_{a=\pi_{\theta}(s)}-\nabla_{a}Q^{\pi_{\theta}}(s,a)|_{a=\pi_{\theta}(s)}\big{\|}^{2}\big{]}$.
A.3 Concept Illustration: Good Approximation in Value does not imply Good Approximation in Gradients
We consider the function fitting problem depicted in Figure 2. As illustrated in the graph, two different functions are fitted given a finite number of sample points, both yielding zero approximation error (measured, say, in the mean-square sense), i.e., they align with the true function at the sample points. In fact, since the number of samples is finite, infinitely many functions can be fitted (exactly). Although the approximation error is identically zero for both functions, the gradients (i.e., the slopes of the lines) do not necessarily align with each other or with the true function, potentially leading to very different gradient directions (at the sample points, in particular).
This observation implies that in sample-based RL methods, regardless of the size of the replay buffer, the possibility of such inconsistency always exists. This is because the $Q$-function approximation of choice is fitted using data sampled from the replay buffer, which is finite. Consequently, the computation of the policy gradient in general becomes inconsistent when the (incompatible, in general) gradient of the fitted $Q$-function is used.
True function: $\frac{1}{25x^{2}+1}$ Function approximation with a polynomial of degree 10
Function approximation with barycentric Lagrange interpolation
Figure 2: Sketch of a function fitting problem containing finite number of samples.
A.4 Proof of Theorem 1
We may write, for every $\theta\in\Theta,\psi,\mu>0$,
$(1-\gamma)\|\hat{\nabla}J(\theta)-\nabla J(\theta)\|$
$=\bigg{\|}\mathbb{E}_{s,\boldsymbol{u}}\bigg{[}\nabla_{\theta}\pi_{\theta}(s)\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\bigg{]}\bigg{|}_{a=\pi_{\theta}(s)}-\mathbb{E}_{s}\big{[}\nabla_{\theta}\pi_{\theta}(s)\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{]}\big{|}_{a=\pi_{\theta}(s)}\bigg{\|}$
$=\bigg{\|}\mathbb{E}_{s}\bigg{[}\nabla_{\theta}\pi_{\theta}(s)\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{]}\bigg{|}_{a=\pi_{\theta}(s)}\bigg{]}\bigg{\|}$
$\leq\mathbb{E}_{s}\bigg{[}\bigg{\|}\nabla_{\theta}\pi_{\theta}(s)\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{]}\bigg{|}_{a=\pi_{\theta}(s)}\bigg{\|}\bigg{]}$
$\leq B\mathbb{E}_{s}\bigg{[}\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{]}\bigg{\|}\bigg{|}_{a=\pi_{\theta}(s)}\bigg{]},$
where we have used Jensen (norms are convex), (matrix) Cauchy-Schwarz, and the assumption that $\sup_{s,\theta}\|\nabla_{\theta}\pi_{\theta}(s)\|\leq B$. In passing, we note that this expression remains valid as $\mu\rightarrow 0$ in which case we obtain the deterministic policy gradient error reading
$(1-\gamma)\|\hat{\nabla}J(\theta)-\nabla J(\theta)\|\leq B\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\psi}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]}.$
We further have, for every $(s,a)\in{\cal S}\times{\cal A}$,
$\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\bigg{]}-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{\|}$
$=\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\bigg{]}-\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)+\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{\|}$
$=\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})}{\mu}\boldsymbol{u}\bigg{]}+\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{\|}$
$\leq\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})}{\mu}\boldsymbol{u}\bigg{]}\bigg{\|}+\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}$
$\leq\dfrac{1}{\mu}\sqrt{\mathbb{E}_{\boldsymbol{u}}[|Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})|^{2}]}\sqrt{\mathbb{E}_{\boldsymbol{u}}[\|\boldsymbol{u}\|^{2}]}+\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}$
$=\dfrac{\sqrt{p}}{\mu}\sqrt{\mathbb{E}_{\boldsymbol{u}}[|Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})|^{2}]}+\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|},$
where we have used Cauchy-Schwarz in the penultimate line. Taking expectations over $s$, we obtain
$\dfrac{(1-\gamma)}{B}\|\hat{\nabla}J(\theta)-\nabla J(\theta)\|$
$\leq\mathbb{E}_{s}\bigg{[}\dfrac{\sqrt{p}}{\mu}\sqrt{\mathbb{E}_{\boldsymbol{u}}[|Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})|^{2}]}+\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\bigg{|}_{a=\pi_{\theta}(s)}\bigg{]}$
$=\dfrac{\sqrt{p}}{\mu}\mathbb{E}_{s}\Big{[}\sqrt{\mathbb{E}_{\boldsymbol{u}}[|Q_{\psi}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})|^{2}]}\Big{]}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]}$
$\leq\dfrac{\sqrt{p}}{\mu}\sqrt{\mathbb{E}_{s,\boldsymbol{u}}[|Q_{\psi}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})|^{2}]}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]}$
$=\dfrac{\sqrt{p}}{\mu}\varepsilon_{\mu,\psi}^{\pi_{\theta}}+\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]},$
where we have used Jensen in the third line (the root function is concave). The rest of the theorem follows trivially from Proposition 1, which is based on the results invoked from Nesterov and Spokoiny [19], and we are done.
A.5 Compatible Policy Gradient with Learnable Exploration Noise
Recent studies have explored learning exploration noise through neural networks. For instance, the closest off-policy RL framework is used by Saglam et al. [22], Saglam and Kozat [21], where the noise term, typically sampled from an isotropic Gaussian, is replaced with the output of a neural network initialized as the policy network. This exploration policy outputs the noise based on the observed state, aiming to maximize uncertainty (i.e., where the loss of the $Q$-network is highest). A natural question that arises is what would happen if we combined such methods with oCPG.
We do not believe it would negatively impact performance because, from the policy perspective, exploratory action noise remains just noise, regardless of the distribution from which it is sampled. Unless the norm of the noise increases significantly, the computed PG would not differ greatly from that computed with Gaussian noise. However, we did not include any learnable exploration method in our experiments, as our focus is solely on studying the pure effects of PG computation, and any additional technique could have confounded our conclusions.
Appendix B Experimental Details
In Section B.1, we provide details of the hyperparameters and network architectures. Additionally, the experimental setup, such as simulation and performance assessment, is examined in Section B.2.
B.1 Hyperparameters and Neural Networks
Architecture
All algorithms use two $Q$-networks following the Clipped Double $Q$-learning algorithm, and one actor network. Each network consists of two hidden layers with 256 units each, using ReLU activations. The $Q$-networks process state-action pairs $(s,a)$ and output a scalar value. The actor network inputs state $s$ and outputs an action $a$ using a tanh activation, scaled to the action range of the environment.
Hyperparameters
The Adam optimizer [10] is used for training the networks, with a learning rate set to $3\times 10^{-4}$. When sampling from the replay buffer, minibatches of 256 samples are employed. We update the target networks using polyak averaging, with a rate of $\tau=0.005$. This process follows the update rule: $\theta^{\prime}\leftarrow 0.995\times\theta^{\prime}+0.005\times\theta$.
Hyperparameter optimization
For SAC, hyperparameter optimization was performed, particularly adjusting the reward scale for the Swimmer task since the original article did not specify this value. We experimented with reward scales of $\{5,10,20\}$ and found that a factor of 5 yielded the best performance in this environment. Apart from this, all algorithms strictly followed the parameter settings and methodologies from their original publications or the latest versions of their code on GitHub. Specifically, SAC adhered to the hyperparameters from its original paper, except for an increase in exploration steps to 25,000. For TD3, we used the version from the author’s GitHub repository2, which introduces slight changes in parameters compared to the paper. Notably, this version increases the start steps to 25,000 and the batch size to 256 across all environments, leading to improved results. The complete hyperparameter settings are listed in Table 2.
Table 2: Hyperparameters used in the experiments.
Hyperparameter Value
Optimizer Adam
Learning rate (all networks) $3\times 10^{-4}$
Learning rate (CPG actor only) $5\times 10^{-5}$
Minibatch size 256
Discount factor $\gamma$ 0.99
Target update rate 0.005
Initial exploration steps 25,000
Exploration noise $\mu$ 0.1
TD3 target policy noise $\sigma$ 0.2
TD3 policy noise clipping $(-0.5,0.5)$
SAC log-standard deviation clipping $(-20,2)$
SAC log constant $10^{-6}$
SAC reward scale (except Humanoid) 5
SAC reward scale (Humanoid) 20
B.2 Experimental Setup
Simulation Environments
All agents are evaluated using continuous control benchmarks from the MuJoCo [^3] https://mujoco.org/ physics engine, accessed through OpenAI Gym [^4] https://www.gymlibrary.ml/ with v3 environments. We maintained the original state-action spaces and reward functions in these environments to ensure reproducibility and fair comparison with existing empirical results. Each environment features a multi-dimensional action space within the range of [-1, 1], except for Humanoid, which has a range of [-0.4, 0.4].
Terminal Transitions
In computing the target $Q$-value, we apply a discount factor of $\gamma=0.99$ for non-terminal transitions, and zero for terminal ones. A transition is considered as terminal only if it ends due to a termination condition, such as a failure or exceeding the time limit or fall of the agent.
Evaluation
Evaluations are performed every 1000 time steps, with each evaluation representing the average reward across 10 episodes. These evaluations use the deterministic policy from oCPG and TD3 without exploration noise, and the deterministic mean action from SAC. To minimize variation due to different seeds, a new environment with a fixed seed (the training seed plus a constant) is used for each evaluation. This approach ensures consistent initial start states for each evaluation.
Visualization of the learning curves
Learning curves, illustrating performance, are plotted as averages from 10 trials. A shaded region around each curve represents the 95% confidence interval across these trials. For enhanced visual clarity, the curves are smoothed using a sliding window averaging over a number evaluations determined by 0.05 of the length of the curves.
B.3 Computational Resources
Our experiments were conducted on a high-performance computing system featuring an AMD Ryzen Threadripper PRO 3995WX 64-Core processor. This system is equipped with 128 logical CPUs operating at a base frequency of 3.31 GHz, with a maximum frequency of 4.31 GHz, supported by 512 GB of RAM. The neural network training leveraged single NVIDIA RTX A6000 GPU with 48 GB of VRAM. This setup provided the necessary computational power to efficiently handle the intensive tasks associated with neural network training and experimentation.
Appendix C Complementary Results
C.1 Numerical Comparison Under Different Statistical Tests
In the main body, we consider Welch’s $t$-test since it accounts for potential differences in variances between the two groups being compared, providing a more robust analysis when these variances may not be equal. Even though our sample sizes are the same, the variance in performance between our method and the baselines could differ, making Welch’s t-test the most appropriate choice to ensure the validity of our statistical significance claims.
To ensure a thorough statistical comparison, we include additional statistical tests in our discussion. Welch’s $t$-test assumes that the data in each group is approximately normally distributed, without equal variances. As a complement, we have chosen the following tests:
  • Student’s $t$-test: Unlike Welch’s test, it compares the means of two independent groups under the assumption that the variances are equal. It is suitable when the distribution of the data is normal, and the variances between groups do not differ significantly.
  • Paired $t$-test: It is used to compare the means of two related groups, such as measurements taken from the same subjects under different conditions. It is ideal for designs where each subject serves as their own control, thereby reducing the impact of variability. Specifically, we aim to reflect the sole difference between oCPG and TD3, which is the PG computation that will serve as the subject’s own control.
  • Wilcoxon rank-sum test: To complement the assumption of Gaussian-distributed sample data, we employ this non-parametric alternative to the independent $t$-test. It compares two independent groups without assuming a normal distribution. It is particularly useful when the data is skewed or contains outliers.
We observe that the statistical significance of the improvement offered by CPG is consistent across different tests, as highlighted in Table 1. This strengthens the validity of our empirical studies and underscores the benefits of our proposed method.
C.2 Learning Curves for Imperfect Environmental Conditions
Setup
The delayed reward setting delays the instantaneous rewards received by the agent by 10 time steps. For noisy rewards, we monitor the range of rewards collected by the agent, specifically the minimum and maximum values. The reward-perturbation noise is then sampled from a normal distribution $\mathcal{N}(0,0.1\cdot(r_{\text{max}}-r_{\text{min}}))$, where $r_{\text{max}}$ and $r_{\text{min}}$ are the maximum and minimum of the rewards in the replay buffer. In the case of sparse rewards, with a random probability of 0.5, the agent either observes the actual instantaneous reward or zero.
TD3       SAC       oCPG
(a) Sparse rewards
(b) Delayed rewards
(c) Noisy rewards
Figure 3: Learning curves for the benchmark MuJoCo environments, averaged over 10 random seeds, under imperfect environment conditions: (a) the agent observes the true reward with a probability of 0.5, otherwise zero; (b) rewards are delayed by 10 time steps; and (c) rewards are perturbed by a zero-mean Gaussian noise with a standard deviation equal to 0.1 of the reward range in the replay buffer. The shaded area represents the 95% confidence interval of the mean performance.