### v2

# Compatible Gradient Approximations for Actor-Critic Algorithms

Abstract
Deterministic policy gradient algorithms are foundational for actor-critic methods in controlling continuous systems, yet they often encounter inaccuracies due to their dependence on the derivative of the critic’s value estimates with respect to input actions. This reliance requires precise action-value gradient computations, a task that proves challenging under function approximation. We introduce an actor-critic algorithm that bypasses the need for such precision by employing a zeroth-order approximation of the action-value gradient through two-point stochastic gradient estimation within the action space. This approach provably and effectively addresses compatibility issues inherent in deterministic policy gradient schemes. Empirical results further demonstrate that our algorithm not only matches but frequently exceeds the performance of current state-of-the-art methods by a substantial extent.
1 Introduction
Reinforcement learning (RL) has established itself as a prominent and effective method for addressing dynamic decision problems [36]. A fundamental strategy in tackling RL problems is via policy gradient (PG) techniques [37], grounded in the hypothesis that actions are selected according to a parameterized distribution. PG methods iteratively update policies via stochastic gradient schemes [37], enabling solutions in complex problems featuring continuous state-action spaces under uncertainty. PG methods are often framed within an actor-critic framework [12], where an actor determines the control policy evaluated by a critic. Thus, the efficacy of the actor’s improvements critically depends on the quality of the feedback provided by the critic. This interdependence raises nontrivial questions on the efficiency of training strategies for both the actor and the critic [5].
Conventionally, the critic is optimized through temporal difference (TD) learning [35] to accurately predict the expected returns led by the actor’s decisions. If the control policy is deterministic, the value estimated by the critic is solely used to compute the action-value gradient during policy optimization [5]. This involves computing the gradient of the critic’s action-value estimates with respect to the actions selected by the policy. More precisely, the deterministic policy gradient (DPG) [33] is determined by chaining the actor’s Jacobian to the action-gradient of the critic. However, TD learning optimizes a critic that learns the action-value, rather than its gradient. Generally, when function approximation is used for the critic in DPG, the resulting PG may not align with the true gradient. In some cases, it may not even constitute an ascent direction [33]. Nevertheless, this failure mode can be circumvented provided that $Q$-function estimates are compatible, meaning a set of conditions necessary for (approximately) following the true gradient; see, e.g., Theorem 3 of Silver et al. [33].
Integration of high-dimensional scenarios requires complex function approximations, such as deep neural networks [14, 8, 9]. However, the use of neural networks violates the classical compatibility results, which (conveniently) calls for critics with linear gradient representations [33, Theorem 3]. Hence, contemporary sample-based algorithms contradict the very foundations of DPG (from the angle of compatibility), thereby paving the way for potential failure scenarios (see Appendix A.2 for an illustration). Indeed, action differentiation of the $Q$-function (estimates), i.e., action-value gradients, leaves modern successors to DPG technically unsound [5].
In this paper, we introduce (off-policy) Compatible Policy Gradient (oCPG), a continuous control actor-critic algorithm that approximates the action-value gradient using only value estimates from the critic. The proposed method embodies a zeroth-order nature by removing the direct computation of the action-value gradient, while provably addressing the compatibility requirement in DPG. This is achieved through evaluating critics at low-dimensional random perturbations within the action space, resulting in (batch) two-point policy gradient approximations whose distance to the true deterministic policy gradient is controllable at will.
The performance of oCPG is assessed on challenging OpenAI Gym continuous control tasks [4], in perfect (less uncertain) and imperfect (higher uncertainty) environmental conditions. Our numerical results demonstrate that oCPG exhibits robust performance and either matches or surpasses (often substantially) the state-of-the-art in terms of stability and mean rewards. For reproducibility purposes, we share our code in the GitHub repository [^1] https://github.com/baturaysaglam/compat-grad-approx.
2 Related Work
Policy gradient techniques are widely used in RL. Various algorithms for estimating policy gradients have been developed; Baxter and Bartlett [3], Metelli et al. [19], Williams [39] focus solely on the policy, while Mnih et al. [20], Schulman et al. [31, 32], Konda and Tsitsiklis [12], Prokhorov and Wunsch [24] also incorporate a value function, known as actor-critic methods. Particularly, algorithms based on the deterministic policy gradient [14, 33] rely on the action-gradient from the critic. In function approximation, the accuracy of the critic is crucial [2]; for example, applying compatibility conditions to the critic helps achieve an unbiased estimate of the policy gradient [33].
The incompatibility of the critic in DPG could be resolved by employing zeroth-order optimization (ZOO), which replaces the action-value gradient computation with an approach that does not require any (first-order) gradient information. These techniques treat the RL problem as a black-box optimization by ignoring the underlying MDP structures and directly searching for the optimal policy without gradient computations. Recent ZOO methods have proven to be competitive on common RL benchmarks [18, 28, 13], benefiting especially from not being constrained to differentiable policies. Despite these advantages, ZOO techniques suffer from high sample complexity and significant variance in gradient updates, due to their disregard for the MDP structure [13]. These issues are further amplified with the increase in perturbation noise scale, potentially escalating dramatically as parameter counts can reach into the millions [6, 13]. Moreover, while parameter space perturbations can induce diverse behaviors viewed as exploration, they are less beneficial in DPG methods, which are typically framed in an off-policy context [33, 14, 8]. Exploration and policy learning are distinct processes in off-policy methods, making parameter-based zeroth-order techniques less effective for enhancing exploration, primarily confined to on-policy and stochastic policies [34]. Contrasting with parameter space-based ZOO techniques, our approach operates in the action space and involves low-dimensional perturbations. Perturbing the action also incorporates Gaussian exploration [8], which has been empirically shown to outperform parameter space exploration in off-policy settings [26, 25]. Lastly, our objective is not merely to devise a pure zeroth-order PG method but to synthesize the strengths of both ZOO and first-order approaches, employing a differentiable policy whose Jacobian is chained to the zeroth-order approximation of the action-value gradient.
In addition to the zeroth-order perspective, GProp [1] and MAGE [5] are the closest known methods to our study that directly focus on optimizing and ensuring accuracy in the action-value gradient computation. GProp modifies traditional TD learning [35] by using an additional neural network that predicts the action-value gradient. MAGE, drawing on recent theoretical [29] and practical [30] insights, trains an additional network to simulate environmental dynamics, employing it as a proxy to sample states for double gradient descent on the $Q$-function. In stark contrast, our method (CPG) simplifies the DPG approach by directly estimating the action-value through a zeroth-order method, employing two-point evaluations of the existing critic network. This eliminates the need to train additional networks, modify TD learning frameworks, or depend on model-based settings. CPG is adaptable to a broader range of problems, bypassing the requirements for double differentiability of the $Q$-function and detailed modeling of transition dynamics. Additionally, our approach seamlessly integrates with any TD learning method and policy learning framework, and requires only a few lines of code to adapt to modern DPG-based algorithms.
3 Technical Preliminaries
We consider a standardized RL problem in which an agent moves through an environment characterized by a continuous compact state space $\mathcal{S}\subset\mathbb{R}^{q}$ and takes actions in a finite-dimensional action space $\mathcal{A}=\mathbb{R}^{p}$. Based on its action selection, the agent receives a reward from a reward function (bounded, for simplicity) $R$, where $R:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$, and observes the next state $s^{\prime}\in\mathcal{S}$. This generic problem is often abstracted by a Markov decision process (MDP) as a tuple $(\mathcal{S},\mathcal{A},P,R,\gamma)$, where $P$ denotes the environment dynamics, that is, the probabilities $P(s^{\prime}|s,a)$ of moving to state $s^{\prime}$ from state $s$ and if action $a$ performed, satisfying the Markov property: $P(s_{t+1}|s_{i},a_{i};\forall i\leq t)=P(s_{t+1}|s_{t},a_{t})$. The value $\gamma\in(0,1)$ is the discount factor that prioritizes future rewards.
The action behavior of the agent is described by a policy $\pi_{\theta}:\mathcal{S}\rightarrow\mathcal{A}$ parameterized by $\theta\in\Theta$, which can be either a deterministic function or represent a probability distribution. Note that we may occasionally omit the parameter subscript. The agent’s objective is to maximize the expected discounted sum of rewards, as quantified by the value function: $V^{\pi_{\theta}}(s)\coloneqq\mathbb{E}[\sum_{t=0}^{\infty}\gamma^{t}R(s_{t},a_{t}\sim\pi_{\theta}(\cdot\mid s_{t}))|s_{0}=s]$. By conditioning on an action, we may also define the action-value function $Q^{\pi_{\theta}}(s,a)\coloneqq\mathbb{E}\left[R(s_{0},a_{0})+\sum_{t=1}^{\infty}\gamma^{t}R(s_{t},a_{t}\sim\pi_{\theta}(\cdot\mid s_{t}))|s_{0}=s,a_{0}=a\right]$. Lastly, let $\rho^{0}$ be the initial state distribution, and let $\rho^{\pi}$ be the $\gamma$-discounted state probability distribution induced by policy $\pi$, defined as $\rho^{\pi}(s^{\prime})=\int_{\mathcal{S}}(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}p_{0}(s)p(s\rightarrow s^{\prime},t,\theta)\mathop{}\!\mathrm{d}s$, where $p(s\rightarrow s^{\prime},t,\theta)$ denotes the probability density (we assume such a density exists for simplicity) at state $s^{\prime}$ after starting from state $s$ and transitioning for $t$ time steps under a policy parameterized by $\theta$.
At the initial state $s$, by selecting the first action according to the policy and advancing the system, the $Q$-function becomes equivalent to the value function. Consequently, the problem is expressed as
$\max_{\theta}\big{\{}J(\theta)\coloneqq\mathbb{E}_{s\sim\rho^{0},a\sim\pi_{\theta}(\cdot|s)}\bigr{[}Q^{\pi_{\theta}}(s,a)\bigr{]}\big{\}}.$
When the search is confined to deterministic policy parameterizations (and we do so hereafter), the objective above reduces to a simpler form, i.e.,
$J(\theta)=\mathbb{E}_{s\sim\rho^{0}}\bigr{[}Q^{\pi_{\theta}}(s,\pi_{\theta}(s))\bigr{]}.$ (1)
Under mild problem regularity conditions, it has been demonstrated by Silver et al. [33] that the gradient of a deterministic policy may be expressed as
$\nabla J(\theta)=\frac{1}{1-\gamma}\int_{\mathcal{S}}\rho^{\pi_{\theta}}(s)\nabla_{\theta}\pi_{\theta}(s)\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{|}_{a=\pi_{\theta}(s)}\mathop{}\!\mathrm{d}s$
$=\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\pi_{\theta}}}\bigr{[}\nabla_{\theta}\pi_{\theta}(s)\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{|}_{a=\pi_{\theta}(s)}\bigr{]},$ (2)
where we use the normalizing term $1/(1-\gamma)$ since $\rho^{\pi_{\theta}}$ is proper. The latter equation is known as the deterministic policy gradient theorem [33], which has been instrumental in the development state-of-the-art policy gradient methods essentially relying on stochastic approximation, such as Deep Deterministic Policy Gradient (DDPG) [14] and Twin Delayed DDPG (TD3) [8].
4 Off-Policy Compatible Policy Gradient
4.1 Problem Statement
Contemporary state-of-the-art (SOTA) first-order PG methods, such as DDPG [14] and TD3 [8], approximate the $Q$-function in (2) using a finite number of samples, which correspond to collected transition tuples. However, accurately approximating function values $Q^{\pi_{\theta}}$ does not necessarily imply accurate gradient approximation $\nabla_{a}Q^{\pi_{\theta}}$, and whether this happens for SOTA methods is unclear and remains an open problem. In fact, for any finite number of samples, it is possible to fit a function with zero approximation error (say, in mean-square sense) but with arbitrarily distinct gradients at each point, unlike the original function. This concept is illustrated with an intuitive example in Appendix A.2. In the context of PG for RL, this issue is very well-known and has been documented in detail in, e.g., [33], among several other influential papers in the literature. The hypothesis that we investigate in this work is that, if such fitting is to be employed whatsoever, then PG should be directly connected to the values of the fitted $Q$-function (using finite samples), rather than the gradients of the fitted $Q$-function.
The accuracy of the approximated PG can still be maintained with a function-approximated Q-function, however rather stringent conditions have to be met [33, Theorem 3]. For completeness, we provide a detailed summary of these compatibility requirements in Appendix A.1. A critical point of Silver et al. [33] for ensuring such compatibility is to conveniently assume that the gradient of the $Q$-function linear in parameters, i.e., $\nabla_{a}Q_{\omega}(s,a)|_{a=\pi_{\theta}(s)}=\nabla_{\theta}\pi_{\theta}(s)^{\top}\omega$ for parameters $\omega$. However, it has been demonstrated that scaling to large and complex state-action spaces inevitably requires the use of more expressive representations, such as deep neural networks [14]. This necessity potentially compromises compatibility and results in $Q$-function gradients that may not accurately follow the true policy gradient [33].
This failure mode has recently been linked to the need to minimize the norm of the action-value gradient of the policy evaluation error, rather than its value, in order to reduce the approximation error of the computed PG [5, Proposition 3.1]. Aligning with this insight, double backpropagation through the TD error may address this issue in scenarios where model-based approaches are feasible [5]. However, model-free RL is often preferred due to its effectiveness in environments with complex or unknown dynamics, offering enhanced flexibility and robustness for direct policy learning via interactions with the environment [36]; this is the path taken herein.
While several studies have pursued the goal of overcoming the incompatibility of the critic [5, 1], our goal herein is to consistently approximate the deterministic policy gradient $\nabla J$ as
^$\hat{\nabla}J(\theta)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\pi}}\big{[}\nabla_{\theta}\pi_{\theta}(s)\hat{\nabla}_{a}Q^{\pi_{\theta}}(s,a))\big{|}_{a=\pi_{\theta}(s)}\big{]},$ (3)
where ^$\hat{\nabla}_{a}Q^{\pi_{\theta}}(s,a)$ represents an appropriately designed approximation of ${\nabla}_{a}Q^{\pi_{\theta}}(s,a)$, resulting in a ^$\hat{\nabla}J$ that is provably compatible, general and implementable in a model-free manner. To the best of our knowledge, no study has yet proposed a gradient surrogate (i.e., ^$\hat{\nabla}J$) that possesses all aforementioned features simultaneously, in a model-free deep RL setting. As we empirically demonstrate later (Section 5), such an approach in fact results in substantial operational benefits over the state-of-the-art.
4.2 Provably Compatible Policy Gradient Approximations
To begin, we temporarily assume that $Q^{\pi_{\theta}}$ is available and, for a smoothing parameter $\mu>0$, we consider the $\mu$-smoothed $Q$-function defined as
$Q^{\pi_{\theta}}_{\mu}(s,a)\coloneqq\mathbb{E}_{\boldsymbol{u}}\big{[}Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})\big{]},\,\,\boldsymbol{u}\sim\mathcal{N}(0,I_{p}),$ (4)
provided that the smoothing is well-defined. Smoothing may be equivalently thought of as enforcing exploration, and the random perturbation $\mu\boldsymbol{u}$ may be thought of as the purely stochastic part of a standard Gaussian policy $\pi_{\theta}^{\mu}(\cdot\mid s)\coloneqq\pi_{\theta}(s)+\mu\boldsymbol{u}$, with $\boldsymbol{u}\sim\mathcal{N}(0,I_{p})$. Here, it is helpful to view such a Gaussian policy as a deterministic policy perturbed by noise in the action space, which is usually of substantially lower dimension as compared with the dimension of the corresponding parameter space [17, 16, 7, 21].
Implementing (Gaussian) smoothing on deterministic actions aligns with approaches used in seminal works such as by Lillicrap et al. [14], Fujimoto et al. [8]. As mentioned above, this technique introduces stochasticity in action selection, which prevents premature convergence to suboptimal policies by promoting exploration of a broader range of state-action pairs [36, 8]. In any case, though, the ultimate goal is to discover a near-optimal deterministic control policy that near-solves (1). In our context, we achieve this by learning a deterministic policy through trajectories generated by a smoothed/perturbed policy, with all performance assessments based on deterministic actions.
Our development will be relying on the following basic result by Nesterov and Spokoiny [23], which is central in the development and analysis of zeroth-order optimization methods, and establishes that the gradient of smoothed functions such as that in (4) may be evaluated in a model-free fashion, through batches of two-point evaluations of the function itself.
Proposition 1 [23].
Let $f:\mathbb{R}^{p}\rightarrow\mathbb{R}$ be a bounded function. For every $\mu>0$, the smoothed function $f_{\mu}(x)\coloneqq\mathbb{E}_{\boldsymbol{u}}\big{[}f(x+\mu\boldsymbol{u})\big{]}$, $\boldsymbol{u}\sim\mathcal{N}(0,I_{p})$ is well-defined, differentiable and its gradient admits the representation
$\nabla f_{\mu}(x)=\mathbb{E}_{\boldsymbol{u}\sim\mathcal{N}(0,I_{n})}\Bigg{[}\frac{f(x+\mu\boldsymbol{u})-f(x)}{\mu}\boldsymbol{u}\Bigg{]},\,\,\forall x\in\mathbb{R}^{p}.$
Further, if $f$ is $G$-smooth (i.e., with $G$-Lipschitz gradients), it holds that
${\sup_{x}}\|\nabla f_{\mu}(x)-\nabla f(x)\|\leq\mu G\sqrt{p}.$
Proposition 1 easily motivates the $Q$-function zeroth-order gradient approximation
${\nabla}_{a}Q_{\mu}^{\pi_{\theta}}(s,a)=\mathbb{E}_{\boldsymbol{u}}\Bigg{[}\frac{Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a)}{\mu}\boldsymbol{u}\Bigg{]},$
resulting in a deterministic policy gradient approximation reading (cf. 3)
^$\hat{\nabla}^{\mu}J(\theta)\coloneqq\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\pi_{\theta}},\boldsymbol{u}\sim\mathcal{N}(0,I_{p})}\Bigg{[}\nabla_{\theta}\pi_{\theta}(s)\times\frac{Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a)}{\mu}\boldsymbol{u}\Bigg{|}_{a=\pi_{\theta}(s)}\Bigg{]}.$
While ^$\hat{\nabla}^{\mu}J$ is a genuine and consistent model-free approximation of the deterministic policy gradient ${\nabla}J$ (through Proposition 1), it requires access to (evaluations of) the $Q$-function itself, the latter being both unknown (in the majority of cases) and also hard, or at least non-trivial to sample.
Approximating the $Q$-function
Following the actor-critic paradigm [12], we approximate the $Q$-function by a parameterized learning representation $Q_{\psi}$ (the critic), for instance, a deep neural network (i.e., $Q$-network), similarly to the control policy $\pi_{\theta}$ (or actor). We denote $Q$-network parameters abstractly with the subscript $\psi$ (different than the smoothing parameter $\mu$). Then, we propose the parameterized deterministic policy gradient approximation
^$\hat{\nabla}^{\mu,\psi}J(\theta)\coloneqq\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\pi_{\theta}},\boldsymbol{u}}\Bigg{[}\nabla_{\theta}\pi_{\theta}(s)\times\frac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\Bigg{|}_{a=\pi_{\theta}(s)}\Bigg{]}.$ (5)
Given (5), a natural question is how the representation $Q_{\psi}$ should be trained, in order for ^$\hat{\nabla}^{\mu,\psi}J$ to achieve a small approximation error as compared with the true policy gradient $\nabla J$. In other words, instead of approximating in value, we would like to choose $\psi$ such that the representation $Q_{\psi}$ is (approximately) compatible to $Q^{\pi_{\theta}}$ (in the standard sense of Silver et al.). To this end, let us first define the perturbation representation error (PRE):
$\varepsilon_{\mu,\psi}^{\pi_{\theta}}\coloneqq\sqrt{\mathbb{E}_{s,\boldsymbol{u}}\big{[}\big{|}Q_{\psi}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})\big{|}^{2}\big{]}},$
where $\theta,\mu$ and $\psi$ are given. We have the following result (see Appendix A.3 for the proof).
Theorem 1 [Compatible Policy Gradient].
Let $B>0$ be a constant such that $\sup_{(s,\theta)\in\mathcal{S}\times\Theta}\|\nabla_{\theta}\pi_{\theta}(s)\|\leq B$. Then the gradient approximation ^$\hat{\nabla}^{\mu,\psi}J$ satisfies, for every $\theta\in\Theta,\mu>0$ and $\psi$,
^$\|\hat{\nabla}^{\mu,\psi}J(\theta)-\nabla J(\theta)\|\leq\dfrac{B}{1-\gamma}\bigg{[}\dfrac{\varepsilon_{\mu,\psi}^{\pi_{\theta}}}{\mu}\sqrt{p}+\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]}\bigg{]}.$
In particular, if $Q^{\pi_{\theta}}(s,\cdot)$ is itself $G$-smooth (uniformly), it follows that, for every $\theta\in\Theta$,
^$\|\hat{\nabla}^{\mu,\psi}J(\theta)-\nabla J(\theta)\|\leq\dfrac{B\sqrt{p}}{1-\gamma}\bigg{[}\dfrac{\varepsilon_{\mu,\psi}^{\pi_{\theta}}}{\mu}+G\mu\bigg{]}.$
The usefulness of the stated result is twofold: On one hand, Theorem 1 quantifies the performance of ^$\hat{\nabla}^{\mu,\psi}J$ in terms of smoothing parameter ($\mu$) and the PRE ($\psi$). On the other hand, Theorem 1 reveals how $\psi$ can be chosen such that the gradient approximation error achieved by adopting ^$\hat{\nabla}^{\mu,\psi}J$ is controlled at will, achieving the desirable policy gradient alignment. To illustrate this, consider a smooth $Q$-function for clarity. Theorem 1 then implies that
$\inf_{\psi}\|\hat{\nabla}^{\mu,\psi}J(\theta)-\nabla J(\theta)\|\leq\dfrac{B\sqrt{p}}{1-\gamma}\bigg{[}\dfrac{{\textstyle\inf_{\psi}\varepsilon_{\mu,\psi}^{\pi_{\theta}}}}{\mu}+G\mu\bigg{]}.$ (6)
We first observe that the smoothing term $\mu$ appears both in the nominator and denominator of the upper bound. For a fixed level of smoothing $\mu>0$ (and at each actor instance indexed by $\theta$), one can optimally train the representation $Q_{\psi}$ to minimize the PRE $\varepsilon_{\mu,\psi}^{\pi_{\theta}}$. Consequently, by selecting a sufficiently expressive class for $Q_{\psi}$, it is theoretically possible to achieve an arbitrarily small gradient approximation error. In particular, one can ideally overfit the critic for each approximate gradient evaluation, and this results in the first term on the right side of (6) disappearing, yielding a total gradient approximation error of the order of $\mathcal{O}(\mu)$. In practice though, this term does not disappear completely. In general, a positive $\mu>0$ must strike a “sweet spot” for good enough compatibility and adequate exploration; this requires sufficiently accurate $Q$-value estimation (in the sense of minimizing the PRE $\varepsilon_{\mu,\psi}^{\pi_{\theta}}$) to maintain balance for a fixed $\mu$, which essentially implements and controls the degree of Gaussian exploration [8]. In any case, minimizing the PRE $\varepsilon_{\mu,\psi}^{\pi_{\theta}}$ over $\psi$ is justified in favor of achieving policy gradient compatibility, in light of Theorem 1. By contrast, standard DPG may suffer from uncontrolled gradient errors due to potential incompatibility (see Section 4.1). Motivated by these facts, we refer to ^$\hat{\nabla}^{\mu,\psi}J$ as a Compatible Policy Gradient (CPG).
In addition to the above, we can also make a couple of further remarks. First, it should be noted that Theorem 1 is general, in it does not assume a specific state distribution (see also proof in the supplement), and considering the on-policy distribution $\rho^{\pi_{\theta}}$ (for each fixed $\theta$) in this section is merely a narrative choice. The arguments leading to and including Theorem 1 apply to any state distribution, e.g., any empirical distribution and in particular that induced from the replay buffer in an off-policy setting (see Section 4.3). Using the notation $\mathbb{E}_{s}[\cdot]$ in the statement of Theorem Theorem 1 highlights that we avoid specifying a particular sampling distribution.
Furthermore, the reader may observe that the “negative” term $Q_{\psi}(s,a)\boldsymbol{u}|_{a=\pi_{\theta}(s)}$ in (5) is of mean zero, and therefore can be eliminated; this is also the case in Proposition 1. Indeed, Theorem 1 holds verbatim for this one-sided gradient surrogate, as well. The reason that the term $Q_{\psi}(s,a)\boldsymbol{u}|_{a=\pi_{\theta}(s)}$ is included in (5) is for ensuring finite variance of the quantities under the expectation $\mathbb{E}_{s,\boldsymbol{u}}$ (formulating a stochastic gradient approximation), essentially playing the role of a baseline estimate. In practice, this extra term improves stability of the resulting gradient approximation, often substantially.
Lastly, it is worth noting that certain convergence results can be obtained using standard methods, such as those involving on-policy actor-critic schemes where the critic is fully optimized before or after each actor update (similar to DDPG/TD3, but for on-policy settings). While this falls outside the scope of our current work, it presents intriguing possibilities for future research.
A final point to emphasize is that although Gaussian exploration noise appears in $\varepsilon_{\mu,\psi}^{\pi_{\theta}}$, the ultimate goal is deterministic policy optimization, as seen in off-policy methods such as DDPG [14, Algorithm 1, lines 8-13] and TD3 [8, Algorithm 1, lines 6-11]. However, both DDPG and TD3 use the gradient of the fitted $Q$-function, which breaks the compatibility, as addressed in Section 4.1. Theorem 1 suggest that (5) can instead be used as a (provably compatible) gradient approximation of choice.
4.3 Off-Policy Deep Reinforcement Learning
To devise an efficient learning framework centered on CPG, we incorporate standard components from existing literature.
Clipped Double $Q$-learning [8]
Based on our discussion above, to exploit ^$\hat{\nabla}^{\mu,\psi}J$ as a gradient surrogate (and thus harness the benefits of Theorem 1), we should learn a $Q_{\psi^{*}}$ such that $\psi^{*}\in\arg\min_{\psi}\varepsilon_{\mu,\psi}^{\pi_{\theta}}$ (iteration-wise for fixed $\theta$ and $\mu$). However, learning such a $Q_{\psi^{*}}$ assumes the feasibility of obtaining unbiased estimates of the $Q$-function at any state-action (perturbation) pair, as well as sampling from the discounted state distribution $\rho^{\pi}$. Specifically for obtaining unbiased $Q$-value estimates in the context of deep RL, one can employ the practical clipped double $Q$-learning algorithm [8] to learn the $Q$-network (in conjunction with an experience replay buffer – see next paragraph), which has been shown to eliminate the estimation bias effectively.
Off-Policy Learning
Our selection of an off-policy learning approach is driven by its potential for high sample efficiency, notably through the use of an experience replay buffer [15]. The agent stores samples as transition tuples $(s,a,R(s,a),s^{\prime})$ in the buffer and periodically samples a minibatch of these transitions for policy and $Q$-network updates. Building upon this, CPG, although theoretically on-policy, is implemented in an off-policy framework. This approach, however, inherently involves a distributional shift, as the samples used may significantly diverge from the current agent’s policy [36]. Despite this, our method draws on the strategy employed by modern off-policy algorithms [33, 14, 8, 9], which simplifies the complexities by approximating off-policy samples as on-policy.
Combining the components introduced, we present our proposed CPG method in an off-policy deep RL setting, implementing stochastic gradient ascent via the implementable gradient approximation
^$\boxed{\hat{\nabla}^{\mu,\psi}J(\theta)\approx\hat{\nabla}_{\beta}^{\mu,\psi}J(\theta)=\frac{1}{1-\gamma}\mathbb{E}_{s\sim\rho^{\beta},\boldsymbol{u}}\Bigg{[}\nabla_{\theta}\pi_{\theta}(s)\frac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\Bigg{|}_{a=\pi_{\theta}(s)}\Bigg{]},}$
where $\beta$ denotes the set of policies that collected the off-policy samples (referring to the replay buffer), which can also be on-policy. Recognizing the limitations of this approximation, we focus on the immediate performance of our algorithm. We note that a detailed analysis of the “off-policyness” (see, e.g., Munos et al., Saglam et al.), as well as the complexity and convergence analysis of this actor-critic setup under CPG, are not the topics of our current scope, and thus deferred for future research.
Algorithm 1 Off-Policy Compatible Policy Gradient (oCPG)
1:Initialize the policy network $\pi_{\theta}$, and critic networks $Q_{\psi_{1}}$, $Q_{\psi_{2}}$ with random parameters $\theta$, $\psi_{1}$, $\psi_{2}$
2:Initialize target networks $\theta^{\prime}\leftarrow\theta$, $\psi^{\prime}_{1}\leftarrow\psi_{1}$, $\psi^{\prime}_{2}\leftarrow\psi_{2}$
3:Initialize the experience replay buffer: $\mathcal{B}=\emptyset$
4:for $t=1$ to $T$ do
5:     Select action with Gaussian exploration noise $a=\pi_{\theta}(s)+\mu\boldsymbol{u}$, where $\boldsymbol{u}\sim\mathcal{N}(0,I_{p})$
6:     Execute action $a$, receive reward $r=R(s,a)$, and observe new state $s^{\prime}$
7:     Store the collected transition tuple in the replay buffer: $\mathcal{B}\leftarrow\mathcal{B}\cup(s,a,r,s^{\prime})$
8:     Sample a minibatch of $N$ transitions from the replay buffer: $(\mathbf{s},\mathbf{a},\mathbf{r},\mathbf{s}^{\prime})\sim\mathcal{B}$
9:     $\tilde{\mathbf{a}}\leftarrow\pi_{\theta^{\prime}}(\mathbf{s}^{\prime})+\epsilon$, where $\{\epsilon_{i}\sim\text{clip}\left(\mathcal{N}(0,\sigma^{2}I_{p}),-c,c\right)\}_{i=1}^{N}$triangleright Clipped Double $Q$-learning
10:     $\mathbf{y}\leftarrow\mathbf{r}+\gamma\text{ min}_{i=1,2}Q_{\psi^{\prime}_{i}}(\mathbf{s}^{\prime},\tilde{\mathbf{a}})$triangleright Clipped Double $Q$-learning
11:     Update critics: $\psi_{j}\leftarrow\text{argmin}_{\psi_{j}}\frac{1}{N}\sum_{i}\left(\mathbf{y}_{i}-Q_{\psi_{j}}(\mathbf{s}_{i},\mathbf{a}_{i})\right)^{2}$triangleright Clipped Double $Q$-learning
12:     if $t$ mod $d$ then $\triangleright$ Delayed policy updates
13:         Update the policy by CPG: ^${\hat{\nabla}_{\beta}^{\mu,\psi}J(\theta)=\dfrac{1}{N}\sum_{i}\nabla_{\theta}\pi_{\theta}(\mathbf{s}_{i})\dfrac{Q_{\psi_{1}}(\mathbf{s}_{i},\mathbf{a}_{i}+\mu\boldsymbol{u}_{i})-Q_{\psi_{1}}(\mathbf{s}_{i},\mathbf{a}_{i})}{\mu}\boldsymbol{u}_{i}\Bigg{|}_{\begin{subarray}{c}\mathbf{a}_{i}=\pi_{\theta}(\mathbf{s}_{i})\\ \boldsymbol{u}_{i}\sim\mathcal{N}(0,I_{p})\end{subarray}}}$
14:         Update target networks: $\theta^{\prime}\leftarrow\tau\cdot\theta+(1-\tau)\cdot\theta^{\prime}$ 15:$\psi_{i}^{\prime}\leftarrow\tau\cdot\psi_{i}+(1-\tau)\cdot\psi_{i}^{\prime}$
16:     end if
17:end for
The pseudocode for the off-policy CPG (oCPG) is provided in Algorithm 1. The inclusion of dual $Q$-networks, target parameters, and delayed policy updates draws from the structure of Clipped Double $Q$-learning. We also consistently optimize the policy with respect to the first $Q$-network to ensure stability.
5 Experiments
We benchmark the performance of oCPG against state-of-the-art PG-based actor-critic algorithms: TD3 [8] and Soft Actor-Critic (SAC) [9]. Since TD3 is a variant of DPG, comparing it with CPG provides a strong test of our proposed PG method against traditional DPG in deep RL. Likewise, as SAC extends the standard stochastic PG algorithm, comparisons with SAC highlight CPG’s effectiveness relative to stochastic PG in the deep RL setting.
5.1 Experimental Setup
Evaluation
Performance is assessed on continuous control tasks from the MuJoCo suite [38] using OpenAI Gym [4]. Each algorithm is trained until convergence in evaluation reward, with evaluations performed every 1,000 steps in a noise-free environment using deterministic actions. Results are averaged over 10 random seeds, capturing variability from the Gym simulator, network initialization, and stochastic processes.
Implementation and Hyperparameters
For simplicity, the implementation of CPG omits the $1-\gamma$ term in the denominator. To ensure a fair comparison with TD3 and isolate the impact of our proposed PG method, we used the same hyperparameters (e.g., learning rate), with the only difference being how the policy gradient is computed.
Selection of $\mu$
In Section 4.2, we highlighted the importance of selecting the smoothing term $\mu$. Initial simulations tested $\mu=\{0.025,0.05,0.075,0.1,0.125\}$. While 0.1 and 0.125 performed similarly, results worsened as $\mu$ decreased. This reflects the nuanced balanced between $\mu$ and the $Q$-function approximation error $\varepsilon_{\mu,\psi}^{\pi_{\theta}}$ in the PG error’s upper bound, i.e., (6). It appears that the $Q$-function approximation could not sufficiently compensate for $\mu<0.1$, suggesting a more expressive $Q$-function approximation. To choose between 0.1 and 0.125, we aligned $\mu$ with the standard deviation of the Gaussian exploration noise used in TD3, which is 0.1. This alignment ensures (i) fair evaluations with TD3, (ii) an adequate exploration, and (iii) a moderate value for smoothing.
A detailed description of the experimental setup and implementation can be found in Appendix B.2.
TD3       SAC       oCPG
Figure 1: Learning curves for benchmark MuJoCo environments, averaged over 10 random seeds. Evaluation reward is calculated as the undiscounted sum of rewards collected by the agent during each evaluation episode. The shaded region indicates the 95% confidence interval of the mean performance.
5.2 Results
The results are presented in Figure 1. Additionally, Table 1 reports the converged rewards, calculated as the mean of the last 50 evaluation rewards, along with several statistical tests to assess the significance of performance differences. We examine these statistical tests in Appendix C.1, highlighting their computational differences and underlying assumptions.
In most environments, oCPG achieves faster convergence to higher final rewards. Although the performance curves are intertwined, making it challenging to visually identify a clear leader, the four statistical tests reported in Table 1 confirm that oCPG is the best-performing algorithm in three out of six environments. In the remaining environments, it is on par with the baselines.
Notice that the only difference between oCPG and TD3 lies in how the PG is computed. All other configurations, including the degree of exploration $\mu$ (which also influences the smoothness in our PG algorithm), remain unchanged. Thus, the statistically significant performance improvement can be attributed to the PG computation. While we do not explicitly claim it, this observation suggests that CPG may produce more accurate PG estimates than DPG.
Environment TD3 SAC oCPG
Ant 4225.080 $\pm$ 835.236 3945.508 $\pm$ 673.329 5284.923 $\pm$ 412.932
HalfCheetah 6744.592 $\pm$ 671.739 6074.230 $\pm$ 405.968 7956.502 $\pm$ 1178.024
Hopper 3620.488 $\pm$ 137.280 3585.355 $\pm$ 186.848 3687.673 $\pm$ 118.093
Humanoid 5289.623 $\pm$ 148.741 5567.597 $\pm$ 141.133 5348.386 $\pm$ 109.030
Swimmer 59.806 $\pm$ 9.477 58.396 $\pm$ 10.990 123.130 $\pm$ 59.253
Walker2d 4842.336 $\pm$ 723.696 4725.314 $\pm$ 492.717 4953.588 $\pm$ 298.321
Table 1: Mean of the last 50 evaluation rewards on benchmark MuJoCo environments, averaged over 10 random seeds, with $\pm$ denoting a 95% confidence interval. The highest performance is marked in boldface. The statistically superior algorithm is highlighted (where applicable), while the highest performances shared by multiple algorithms are indicated. Welch’s test, Student’s t-test, paired t-test, and the Wilcoxon rank-sum test are conducted at a significance level of 0.05, all yielding consistent results across methods.
We observe that oCPG converges more slowly than TD3 in the Humanoid environment, where the combined state-action dimension is 393. In our experiments, the $Q$-networks have only 256 hidden units per layer, and neural networks are sensitive to input noise, which becomes more noticeable as the input size surpasses the network’s capacity. This issue arises due to the smoothing step on line 13 in Algorithm 1; however, using larger networks could mitigate this limitation. Despite this, oCPG achieves the highest converged rewards in five of six environments. While its performance in the Swimmer task is statistically significant, the rewards show high variance because some seeds converge to suboptimal levels, leading to wide confidence intervals.
Moreover, discrepancies in reported reward levels compared to the TD3 and SAC papers may result from differences in environmental stochasticity and the MuJoCo simulator version. Nevertheless, consistent performance across multiple random seeds provides a reliable basis for evaluating our method. This approach aligns with established deep RL experimentation standards [10], which prioritize comparative analysis over absolute performance.
Finally, the smoothing parameter $\mu$ is linked to exploration levels, eliminating the need for environment-specific tuning. Our experiments show that a fixed value of $\mu=0.1$ generalizes well across diverse environments, ensuring effective policy gradient approximation while preserving exploratory behavior.
Imperfect Environmental Conditions
To evaluate the performance of our method under uncertainty, we present results with perturbed rewards in Appendix C. Convergence is noticeably affected in environments with altered rewards, as uncertainty distorts the agent’s received rewards, directly influencing the $Q$-value used by CPG for policy gradient computation. Although the baselines also depend on $Q$-value estimates for their gradients, reward modifications have minimal effect on their final performance. This demonstrates that the precision of value gradient estimates does not necessarily correlate with the accuracy of the value estimates themselves, as discussed in Appendix A.2. Despite these challenges, oCPG achieves higher reward levels over time.
6 Conclusions and Future Work
We introduced Compatible Policy Gradient (CPG), a policy gradient technique for actor-critic algorithms that employs a zeroth-order approximation method to estimate the action-value gradient in deterministic policy gradient (DPG) algorithms. By employing stochastic two-point gradient estimation with low-dimensional perturbations in action space, our approach eliminates the need for explicit action-value gradient computation, addressing compatibility issues commonly faced in traditional DPG methods with deep function approximation. Empirical results on various continuous control tasks demonstrate that CPG outperforms state-of-the-art methods.
Our findings highlight two key implications. First, they reveal the potential of zeroth-order (gradient-free) methods to overcome challenges related to accurate derivative computations in policy optimization. Second, the strong performance of CPG in complex continuous control environments suggests that similar approaches could prove effective in other reinforcement learning domains characterized by uncertain dynamics or imperfect system models.
Future Work
An off-policy correction mechanism for CPG could be developed by leveraging the probability ratio between the agent’s current policy and its behavioral policy. Additionally, while we did not investigate convergence results for on-policy actor-critic methods with fully optimized critics that achieve near-zero approximation error, this remains a compelling direction for future research.
Appendix A Supporting Details for Section 4
A.1 Compatibility Requirements in Deterministic Policy Gradients
Here, for completeness we recall the classical result of Silver et al. [33] (specifically Theorem 3 in Silver et al. [33]), which ensures exact gradient compatibility, however at the expense of restrictive conditions on the form of the adopted $Q$-function approximation (in particular, its gradient must be of linear form).
Theorem 2 Compatible Function Approximation.
A function approximator $Q_{\omega}(s,a)$ is compatible with a deterministic policy $\pi_{\theta}(s)$, $\nabla J(\theta)=\mathbb{E}_{s\sim\rho^{\pi_{\theta}}}\bigr{[}\nabla_{\theta}\pi_{\theta}(s)\nabla_{a}Q_{\omega}(s,a)\big{|}_{a=\pi_{\theta}(s)}\bigr{]}$, if
  1. 1.
    $\nabla_{a}Q_{\omega}(s,a)\big{|}_{a=\pi_{\theta}(s)}=\nabla_{\theta}\pi_{\theta}(s)^{\top}\omega$ and
  2. 2.
    $\omega$ minimizes the gradient mean-squared error $\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\omega}(s,a)|_{a=\pi_{\theta}(s)}-\nabla_{a}Q^{\pi_{\theta}}(s,a)|_{a=\pi_{\theta}(s)}\big{\|}^{2}\big{]}$.
A.2 Concept Illustration: Accurate Value Approximation does not Guarantee Accurate Gradient Approximation
We consider the function fitting problem depicted in Figure 2. As illustrated in the graph, two different functions are fitted given a finite number of sample points, both yielding zero approximation error (measured, say, in the mean-square sense), i.e., they align with the true function at the sample points. In fact, since the number of samples is finite, infinitely many functions can be fitted (exactly). Although the approximation error is identically zero for both functions, the gradients (i.e., the slopes of the lines) do not necessarily align with each other or with the true function, potentially leading to very different gradient directions (at the sample points, in particular).
This observation implies that in sample-based RL methods, regardless of the size of the replay buffer, the possibility of such inconsistency always exists. This is because the $Q$-function approximation of choice is fitted using data sampled from the replay buffer, which is finite. Consequently, the computation of the policy gradient in general becomes inconsistent when the (incompatible, in general) gradient of the fitted $Q$-function is used.
True function: $\frac{1}{25x^{2}+1}$ Function approximation with a polynomial of degree 10
Function approximation with barycentric Lagrange interpolation
Figure 2: Sketch of a function fitting problem containing finite number of samples.
A.3 Proof of Theorem 1
We may write, for every $\theta\in\Theta,\psi,\mu>0$,
$(1-\gamma)\|\hat{\nabla}J(\theta)-\nabla J(\theta)\|$
$=\bigg{\|}\mathbb{E}_{s,\boldsymbol{u}}\bigg{[}\nabla_{\theta}\pi_{\theta}(s)\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\bigg{]}\bigg{|}_{a=\pi_{\theta}(s)}-\mathbb{E}_{s}\big{[}\nabla_{\theta}\pi_{\theta}(s)\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{]}\big{|}_{a=\pi_{\theta}(s)}\bigg{\|}$
$=\bigg{\|}\mathbb{E}_{s}\bigg{[}\nabla_{\theta}\pi_{\theta}(s)\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{]}\bigg{|}_{a=\pi_{\theta}(s)}\bigg{]}\bigg{\|}$
$\leq\mathbb{E}_{s}\bigg{[}\bigg{\|}\nabla_{\theta}\pi_{\theta}(s)\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{]}\bigg{|}_{a=\pi_{\theta}(s)}\bigg{\|}\bigg{]}$
$\leq B\mathbb{E}_{s}\bigg{[}\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{]}\bigg{\|}\bigg{|}_{a=\pi_{\theta}(s)}\bigg{]},$
where we have used Jensen (norms are convex), (matrix) Cauchy-Schwarz, and the assumption that $\sup_{s,\theta}\|\nabla_{\theta}\pi_{\theta}(s)\|\leq B$. In passing, we note that a non-standard (i.e., transposed) definition for the Jacobian $\nabla_{\theta}\pi_{\theta}(s)$ was used when applying the chain rule. Additionally, the latter expression remains valid as $\mu\rightarrow 0$ in which case we obtain the deterministic policy gradient error reading
$(1-\gamma)\|\hat{\nabla}J(\theta)-\nabla J(\theta)\|\leq B\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\psi}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]}.$
We further have, for every $(s,a)\in{\cal S}\times{\cal A}$,
$\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\bigg{]}-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{\|}$
$=\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q_{\psi}(s,a)}{\mu}\boldsymbol{u}\bigg{]}-\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)+\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{\|}$
$=\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})}{\mu}\boldsymbol{u}\bigg{]}+\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\bigg{\|}$
$\leq\bigg{\|}\mathbb{E}_{\boldsymbol{u}}\bigg{[}\dfrac{Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})}{\mu}\boldsymbol{u}\bigg{]}\bigg{\|}+\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}$
$\leq\dfrac{1}{\mu}\sqrt{\mathbb{E}_{\boldsymbol{u}}[|Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})|^{2}]}\sqrt{\mathbb{E}_{\boldsymbol{u}}[\|\boldsymbol{u}\|^{2}]}+\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}$
$=\dfrac{\sqrt{p}}{\mu}\sqrt{\mathbb{E}_{\boldsymbol{u}}[|Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})|^{2}]}+\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|},$
where we have used Cauchy-Schwarz in the penultimate line. Taking expectations over $s$, we obtain
$\dfrac{(1-\gamma)}{B}\|\hat{\nabla}J(\theta)-\nabla J(\theta)\|$
$\leq\mathbb{E}_{s}\bigg{[}\dfrac{\sqrt{p}}{\mu}\sqrt{\mathbb{E}_{\boldsymbol{u}}[|Q_{\psi}(s,a+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,a+\mu\boldsymbol{u})|^{2}]}+\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\bigg{|}_{a=\pi_{\theta}(s)}\bigg{]}$
$=\dfrac{\sqrt{p}}{\mu}\mathbb{E}_{s}\Big{[}\sqrt{\mathbb{E}_{\boldsymbol{u}}[|Q_{\psi}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})|^{2}]}\Big{]}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]}$
$\leq\dfrac{\sqrt{p}}{\mu}\sqrt{\mathbb{E}_{s,\boldsymbol{u}}[|Q_{\psi}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})-Q^{\pi_{\theta}}(s,\pi_{\theta}(s)+\mu\boldsymbol{u})|^{2}]}$
$\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad\quad+\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]}$
$=\dfrac{\sqrt{p}}{\mu}\varepsilon_{\mu,\psi}^{\pi_{\theta}}+\mathbb{E}_{s}\big{[}\big{\|}\nabla_{a}Q_{\mu}^{\pi_{\theta}}(s,a)-\nabla_{a}Q^{\pi_{\theta}}(s,a)\big{\|}\big{|}_{a=\pi_{\theta}(s)}\big{]},$
where we have used Jensen in the third line (the root function is concave). The rest of the theorem follows trivially from Proposition 1, which is based on the results invoked from Nesterov and Spokoiny [23], and we are done.
A.4 Compatible Policy Gradient with Learnable Exploration Noise
Recent studies have explored learning exploration noise through neural networks. For instance, the closest off-policy RL framework is used by Saglam et al. [26], Saglam and Kozat [25], where the noise term, typically sampled from an isotropic Gaussian, is replaced with the output of a neural network initialized as the policy network. This exploration policy outputs the noise based on the observed state, aiming to maximize uncertainty (i.e., where the loss of the $Q$-network is highest). A natural question that arises is what would happen if we combined such methods with oCPG.
We do not believe it would negatively impact performance because, from the policy perspective, exploratory action noise remains just noise, regardless of the distribution from which it is sampled. Unless the norm of the noise increases significantly, the computed PG would not differ greatly from that computed with Gaussian noise. However, we did not include any learnable exploration method in our experiments, as our focus is solely on studying the pure effects of PG computation, and any additional technique could have confounded our conclusions.
Appendix B Experimental Details
In Section B.1, we provide details of the hyperparameters and network architectures. Additionally, the experimental setup, such as simulation and performance assessment, is examined in Section B.2.
B.1 Hyperparameters and Neural Networks
Architecture
All algorithms use two $Q$-networks following the Clipped Double $Q$-learning algorithm, and one actor network. Each network consists of two hidden layers with 256 units each, using ReLU activations. The $Q$-networks process state-action pairs $(s,a)$ and output a scalar value. The actor network inputs state $s$ and outputs an action $a$ using a tanh activation, scaled to the action range of the environment.
Hyperparameters
The Adam optimizer [11] is used for training the networks, with a learning rate set to $3\times 10^{-4}$. When sampling from the replay buffer, minibatches of 256 samples are employed. We update the target networks using polyak averaging, with a rate of $\tau=0.005$. This process follows the update rule: $\theta^{\prime}\leftarrow 0.995\times\theta^{\prime}+0.005\times\theta$.
Hyperparameter Optimization
For SAC, hyperparameter optimization was performed, particularly for the Swimmer task, as the original paper did not specify the reward scale. We tested reward scales of $\{5,10,20\}$ and found that a factor of 5 delivered the best performance in this environment. Beyond this, all algorithms strictly followed the parameter settings from their papers or the latest GitHub versions of their code. Specifically, SAC followed the hyperparameters from its original paper, with the exception of increasing exploration steps (i.e., a purely exploratory policy with no learning) to 25,000. For TD3, we used the author’s GitHub repository [^2] https://github.com/sfujim/TD3, which introduced slight parameter changes compared to the paper. Notably, this version increased the start steps to 25,000 and the batch size to 256 across all environments, resulting in improved performance. The full hyperparameter settings are provided in Table 2.
Hyperparameter Value
Optimizer Adam
Learning rate (all networks) $3\times 10^{-4}$
Learning rate (CPG actor only) $5\times 10^{-5}$
Minibatch size 256
Discount factor $\gamma$ 0.99
Target update rate 0.005
Initial exploration steps 25,000
Exploration noise $\mu$ 0.1
TD3 target policy noise $\sigma$ 0.2
TD3 policy noise clipping $(-0.5,0.5)$
SAC log-standard deviation clipping $(-20,2)$
SAC log constant $10^{-6}$
SAC reward scale (except Humanoid) 5
SAC reward scale (Humanoid) 20
Table 2: Hyperparameters used in the experiments, applied to all methods unless specified otherwise.
B.2 Experimental Setup
Simulation Environments
All agents are evaluated using continuous control benchmarks from the MuJoCo [^3] https://mujoco.org/ physics engine, accessed through OpenAI Gym [^4] https://www.gymlibrary.ml/ with v3 environments. We maintained the original state-action spaces and reward functions in these environments to ensure reproducibility and fair comparison with existing empirical results. Each environment features a multi-dimensional action space within the range of [-1, 1], except for Humanoid, which has a range of [-0.4, 0.4].
Terminal Transitions
In computing the target $Q$-value, we apply a discount factor of $\gamma=0.99$ for non-terminal transitions, and zero for terminal ones. A transition is considered as terminal only if it ends due to a termination condition, such as a failure or exceeding the time limit or fall of the agent.
Evaluation
Evaluations are performed every 1000 time steps, with each evaluation representing the average reward across 10 episodes. These evaluations use the deterministic policy from oCPG and TD3 without exploration noise, and the deterministic mean action from SAC. To minimize variation due to different seeds, a new environment with a fixed seed (the training seed plus a constant) is used for each evaluation. This approach ensures consistent initial start states for each evaluation.
Visualization of the Learning Curves
Learning curves, illustrating performance, are plotted as averages from 10 trials. A shaded region around each curve represents the 95% confidence interval across these trials. For enhanced visual clarity, the curves are smoothed using a sliding window averaging over a number evaluations determined by 0.05 of the length of the curves.
B.3 Computational Resources
Our experiments were conducted on a high-performance computing system with an AMD Ryzen Threadripper PRO 3995WX processor (64 cores) and 512 GB of RAM. Neural network training utilized a single NVIDIA RTX A6000 GPU with 48 GB of VRAM, providing sufficient computational resources for our tasks.
Appendix C Complementary Results
C.1 Numerical Comparison Under Different Statistical Tests
In the main body, we consider four different statistical tests. To ensure a comprehensive statistical comparison, we provide brief descriptions of each method and their approach to conducting statistical analysis.
  • Welch’s $t$-test: It accounts for potential differences in variances between the two groups being compared, providing a more robust analysis when variances are unequal. Although our sample sizes are identical, differences in performance variance between our method and the baselines could exist, making Welch’s $t$-test the most appropriate choice to ensure the validity of our statistical significance claims.
  • Student’s $t$-test: Unlike Welch’s test, it compares the means of two independent groups under the assumption that the variances are equal. It is suitable when the distribution of the data is normal, and the variances between groups do not differ significantly.
  • Paired $t$-test: It is used to compare the means of two related groups, such as measurements taken from the same subjects under different conditions. It is ideal for designs where each subject serves as their own control, thereby reducing the impact of variability. Specifically, we aim to reflect the sole difference between oCPG and TD3, which is the PG computation that will serve as the subject’s own control.
  • Wilcoxon rank-sum test: To complement the assumption of Gaussian-distributed sample data, we employ this non-parametric alternative to the independent $t$-test. It compares two independent groups without assuming a normal distribution. It is particularly useful when the data is skewed or contains outliers.
We observe that the statistical significance of the improvement offered by CPG is consistent across different tests, as highlighted in Table 1. This strengthens the validity of our empirical studies and underscores the benefits of our proposed method.
C.2 Learning Curves for Imperfect Environmental Conditions
Uncertainty in the Environments
To evaluate the robustness of oCPG, we simulate imperfect environmental conditions by introducing uncertainty. This includes considering scenarios with noisy, sparse, and delayed rewards. In the noisy reward scenario, each reward is subject to perturbation with random noise. For the sparse reward setup, the agent observes the reward with a probability less than 1; otherwise, it receives a reward of zero. In the delayed setting, the agent observes the reward after a fixed number of time steps.
Setup
The delayed reward setting delays the instantaneous rewards received by the agent by 10 time steps. For noisy rewards, we monitor the range of rewards collected by the agent, specifically the minimum and maximum values. The reward-perturbation noise is then sampled from a normal distribution $\mathcal{N}(0,0.1\cdot(r_{\text{max}}-r_{\text{min}}))$, where $r_{\text{max}}$ and $r_{\text{min}}$ are the maximum and minimum of the rewards in the replay buffer. In the case of sparse rewards, with a random probability of 0.5, the agent either observes the actual instantaneous reward or zero.
TD3       SAC       oCPG
(a) Sparse rewards
(b) Delayed rewards
(c) Noisy rewards
Figure 3: Learning curves for the benchmark MuJoCo environments, averaged over 10 random seeds, under imperfect environment conditions: (a) the agent observes the true reward with a probability of 0.5, otherwise zero; (b) rewards are delayed by 10 time steps; and (c) rewards are perturbed by a zero-mean Gaussian noise with a standard deviation equal to 0.1 of the reward range in the replay buffer. The shaded area represents the 95% confidence interval of the mean performance.